### v2

# Peri-LN: Revisiting Layer Normalization in the Transformer Architecture

Abstract
Designing Transformer architectures with the optimal layer normalization (LN) strategy that ensures large-scale training stability and expedite convergence has remained elusive, even in this era of large language models (LLMs). To this end, we present a comprehensive analytical foundation for understanding how different LN strategies influence training dynamics in large-scale Transformer training. Until recently, Pre-LN and Post-LN have long dominated standard practices despite their limitations in large-scale training. However, several open-source large-scale models have recently begun silently adopting a third strategy without much explanation. This strategy places layer normalization (LN) peripherally around sublayers, a design we term Peri-LN. While Peri-LN has demonstrated promising empirical performance, its precise mechanisms and benefits remain almost unexplored. Our in-depth analysis shows that Peri-LN strikes an ideal balance in variance growth—unlike Pre-LN and Post-LN, which are prone to vanishing gradients and “massive activations.” To validate our theoretical insight, we conduct large-scale experiments on Transformers up to $3.2$B parameters, showing that Peri-LN consistently achieves more balanced variance growth, steadier gradient flow, and convergence stability. Our results suggest that Peri-LN warrants broader consideration for large-scale Transformer architectures, providing renewed insights into the optimal placement and application of LN.
normalization, transformers, pre-training, large language models, training stability
1 Introduction
Building on a rapidly expanding lineage of Transformer-based large language models, open-source models have shown remarkable impact (Hoffmann et al., 2022; Dubey et al., 2024; Guo et al., 2025). As the demand for larger and more powerful models grows, various training stabilization techniques have been introduced (Yang et al., 2022; Zhai et al., 2023; Loshchilov et al., 2024). Among these, the choice of where and how to apply layer normalization (LN: LayerNorm or RMSNorm; Ba et al., 2016; Zhang & Sennrich, 2019) critically influences model convergence (Xiong et al., 2020; Kedia et al., 2024; Wortsman et al., 2024; Chung et al., 2024). However, their immense computational requirements have restricted deeper exploration of the underlying Transformer structure. Are we truly employing the optimal LN placement? In practice, fully revealing the results of massive resource investments can be challenging (Rivière et al., 2024). Despite its importance, there is still no consensus on a single best LN placement strategy.
Two prominent LN placements have been widely explored. Post-LN (Vaswani et al., 2017) normalizes the hidden state after adding the sub-layer output to the residual stream (that is, $\mathrm{Norm}(x+\mathrm{Module}(x))$ where $x$ is input hidden state. $\mathrm{Norm}$ is LN). This helps constrain the variance of hidden states but may inadvertently weaken gradient signals, particularly in deeper models (Kedia et al., 2024). Pre-LN (Dubey et al., 2024), by contrast, normalizes before passing the hidden state to the sub-layer (that is, $x+\mathrm{Module}(\mathrm{Norm}(x))$). While this can enhance gradient propagation, it also admits so-called “massive activations,” where hidden states grow exponentially across layers (Sun et al., 2024; Wortsman et al., 2024; Zhai et al., 2023).
Previous studies on deep convolutional neural networks (CNNs) have analyzed the impact of batch normalization on variance changes during the initialization stage of ResNet architectures, demonstrating its relationship to model performance (De & Smith, 2020). They noted that, in models without normalization, hidden activation growth at initialization can be exponential, leading to poor performance and stability. In contrast, in pre-normalized CNNs, the variance of hidden activations was shown to increase linearly as model depth grows. In the same vein, Kedia et al. (2024) reported that, for Transformer architectures as well, the variance in the forward propagation of Transformer-based language models at initialization increases linearly with depth. However, in the context of Transformer architectures, we observed that this variance growth at initialization does not persist as training progresses as shown in Figure 1. Section 3 and 4 provide a more detailed discussion of these hidden-state growth patterns.
Beyond these two common strategies, Post-LN and Pre-LN, a third LN placement has quietly emerged in large-scale open-source models: applying LN around the sub-layer, i.e., on both its input and output. Although recent open-source models (Rivière et al., 2024; OLMo et al., 2024) have quietly adopted such designs and demonstrated promising performance on a large scale, these efforts often appeared isolated, lacking a conceptual unifying framework or a thorough investigation into their benefits. In this paper, we coin the term Peri-LN111“Peri-” means “around,” reflecting that LN encapsulates the entire sub-layer. to unify these scattered approaches and highlight an underexplored avenue for stabilizing large-scale Transformer training. While prior work offers encouraging empirical results, there is still a lack of in-depth analysis explaining why and how Peri-LN can help mitigate issues that frequently arise with Post-LN and Pre-LN. By probing the forward and backward dynamics of Peri-LN, we provide fresh insights into where and when it may offer advantages over more widely adopted LN placements.
Accordingly, this paper revisits LN placement in Transformers from both analytical and empirical perspectives. In particular, we:
  1. 1.
    Present an in-depth empirical analysis of Post-LN and Pre-LN in large-scale Transformers, examining how variance and gradient properties evolve beyond initialization. By studying forward and backward propagation statistics, we clarify why Post-LN may suffer vanishing gradients and why Pre-LN can exhibit exponential growth of hidden state.
  2. 2.
    Investigate Peri-LN to understand how normalizing both the inputs and outputs of each module moderates hidden-state behavior during forward and backward propagation, providing a systematic perspective on this underexplored alternative. We show that Peri-LN not only improves gradient stability and final loss but also plays a critical role in reducing hidden-state redundancy, as demonstrated by our ablation studies.
  3. 3.
    Provide quantitative evidence on how large activation influences training stability, benchmark performance, and redundancy in hidden states. Drawing on theoretical bounds for gradient propagation as well as extensive experiments, we demonstrate that Peri-LN can mitigate excessive variance growth and gradient instability more effectively than commonly used alternatives.
Figure 1: Illustration of hidden-state variance across different model depths and training iterations. From the initialization stage up to the point where 6.3 billion tokens were trained, we observed the variance growth of hidden states for Pre-LN and Post-LN architectures. The analysis was conducted using a $1.5$B-parameter model, and consistent trends were observed across models of different sizes. Detailed settings and more results are in Section 4.4.2.
2 Background and Motivation
The analysis of activation variance at model initialization has long been central to understanding normalization layers and enhancing stability in convolutional neural networks (CNNs) (De & Smith, 2020; He et al., 2016; Brock et al., 2021a). De & Smith (2020) showed that batch normalization in residual blocks can bias networks toward the identity function, thereby stabilizing gradients and improving overall training dynamics.
Similar investigations have emerged for Transformer architectures, examining how variance propagates and how gradients behave in both post-layer normalization (Post-LN) (Vaswani et al., 2017) and pre-layer normalization (Pre-LN) (Dubey et al., 2024) configurations (Xiong et al., 2020; Kedia et al., 2024; Wortsman et al., 2024; Li et al., 2024b). Early work comparing Post- and Pre-LN primarily focused on gradient scales and loss behavior. Xiong et al. (2020) observed that Pre-LN architectures tend to exhibit more stable gradients, but can still encounter issues such as gradient spikes and divergence, especially in deeper models or large-scale pre-training scenarios (Zhai et al., 2023; Wortsman et al., 2024; Fishman et al., 2024; Chung et al., 2024).
Among these challenges, the phenomenon of “massive activations” has attracted particular attention (Dettmers et al., 2022; Yu et al., 2024; Fishman et al., 2024). Sun et al. (2024) identified that in Pre-LN architectures, large spikes in activation magnitude can persist across layers due to residual connections. These massive activations act as fixed biases, potentially narrowing the model’s focus to certain tokens and may influence generalization. However, the underlying mechanisms behind these large values—and their exact impact on the training process—remain not yet well understood.
Analytical work has provided theoretical frameworks to explain phenomena like gradient explosion and vanishing in Transformers. For instance, Kedia et al. (2024) introduced a signal propagation theory that details how activation variance and gradient instability can evolve with depth, identifying critical factors that impair stability and performance. Recent studies have discussed how Pre-LN architectures can allow large values from Attention or MLP modules to flow unimpeded through residual connections (Csordás et al., 2024; Fishman et al., 2024; Zhai et al., 2023; Wortsman et al., 2024), but the precise impact of this behavior on large-scale training remains insufficiently explored.
These observations underscore the ongoing need to clarify how activation dynamics, normalization strategies, and architectural choices interact, especially in large-scale models. In response, this work aims to deepen our understanding of activation evolution during Transformer training under different normalization architectures, focusing on the role of massive activations and their effects on overall stability and performance.
We defer an extended discussion of the related literature to Appendix A, owing to space limitations.
3 Comparative Analysis
In this section, we discuss how different placements of layer normalization (LN  [^2] Unless stated otherwise, LN refers to both LayerNorm and RMSNorm.) in Transformer architecture affect both training stability and the statistics of hidden states (activations  [^3] We use “hidden state” and “activation” interchangeably.).
3.1 Post- & Pre-Normalization in Transformers
Post-LN.
The Post-Layer Normalization (Post-LN) (Vaswani et al., 2017) scheme, normalization is applied after summing the module’s output and residual input:
$y_{l}=\mathrm{Norm}\bigl{(}x_{l}+\mathrm{Module}(x_{l})\bigr{)},$ (1)
where $x_{l}$ is the input hidden state of $l$-th layer, $y_{l}$ is the output hidden state of $l$-th layer, and $\mathrm{Module}$ denotes Attention or Multi-Layer Perceptron (MLP) module in the Transformer sub-layer. $\mathrm{Norm}$ denotes normalization layers such as RMSNorm or LayerNorm. It is known that by stabilizing the activation variance at a constant scale, Post-LN prevents activations from growing. However, several evidence (Xiong et al., 2020; Kedia et al., 2024) suggest that Post-LN can degrade gradient flow in deeper networks, leading to vanishing gradients and slower convergence.
Pre-LN.
The Pre-Layer Normalization (Pre-LN) (Dubey et al., 2024) scheme, normalization is applied to the module’s input before processing:
$y_{l}=x_{l}+\mathrm{Module}\bigl{(}\mathrm{Norm}(x_{l})\bigr{)}.$ (2)
As for Llama $3$ architecture, a final LN is applied to the network output. Pre-LN improves gradient flow during backpropagation, stabilizing early training (Xiong et al., 2020). Nonetheless, in large-scale Transformers, even Pre-LN architectures are not immune to instability during training (Wortsman et al., 2024; Zhai et al., 2023). As shown in Figure 2, unlike Post-LN—which places LN at position $C$—Pre-LN, which places LN only at position $A$, can lead to a “highway” structure that is continuously maintained throughout the entire model if the module produces an output with a large magnitude. This phenomenon might be related to the “massive activations” observed in trained models (Sun et al., 2024; Fishman et al., 2024).
A B C
Post-LN × ×
Pre-LN × ×
Peri-LN ×
Figure 2: Placement of normalization in Transformer sub-layer.
Table 1: Intuitive comparison of normalization strategies.
Strategy Variance Growth Gradient Stability
Post-LN Mostly constant Potential for vanishing
Pre-LN Exponential in depth Potential for explosion
[…]
3.2 Variance Behavior from Initialization to Training
As discussed by Xiong et al. (2020) and Kedia et al. (2024), Transformer models at initialization exhibit near-constant hidden-state variance under Post-LN and linearly increasing variance under Pre-LN. Most of the previous studies have concentrated on this early-stage behavior. However, Recent studies have also reported large output magnitudes in both the pre-trained attention and MLP modules (Dehghani et al., 2023; Wortsman et al., 2024; Fishman et al., 2024). To bridge the gap from initialization to the fully trained stage, we extend our empirical observations in Figure 1 beyond initial conditions by tracking how these variance trends evolve at intermediate points in training.
We find that Post-LN maintains a roughly constant variance, which helps avert exploding activations. Yet as models grow deeper and training proceeds, consistently normalizing $x_{l}+\mathrm{Module}(x_{l})$ can weaken gradient flow, occasionally causing partial vanishing gradients and slower convergence. In contrast, Pre-LN normalizes $x_{l}$ before the module but leaves the module output unnormalized, allowing hidden-state variance to accumulate exponentially once parameter updates amplify the input. Although Pre-LN preserves gradients more effectively in earlier stages, this exponential growth in variance can lead to “massive activations” (Sun et al., 2024), risking numeric overflow and destabilizing large-scale training. We reconfirm this in Section 4.
Takeaways.
  • Keeping the Highway Clean: Post-LN’s Potential for Gradient Vanishing and Slow Convergence. When layer normalization is placed directly on the main path (Placement $C$ in Figure 2), it can cause gradient vanishing and introduce fluctuations in the gradient scale, potentially leading to instability.
  • Maintaining a Stable Highway: Pre-LN May Not Suffice for Training Stability. Pre-LN does not normalize the main path of the hidden states, thereby avoiding the issues that Post-LN encounters. Nevertheless, a structural characteristic of Pre-LN is that any large values arising in the attention or MLP modules persist through the residual identity path. In particular, as shown in Figure 1, the exponentially growing magnitude and variance of the hidden states in the forward path may lead to numerical instability and imbalance during training.
Recent open-sourced Transformer architectures have adopted normalization layers in unconventional placements. Models like Gemma2222 and OLMo2222 utilize normalization layers at the module output (Output-LN), but the benefits of these techniques remain unclear (Rivière et al., 2024; OLMo et al., 2024). To investigate the impact of adding an Output-LN, we explore the peri-layer normalization architecture.
3.3 Placing Module Output Normalization
Peri-LN.
The Peri-Layer Normalization (Peri-LN) applies LN twice within each layer—before and after the module—and further normalizes the input and final output embeddings. Formally, for the hidden state $x_{l}$ at layer $l$:
  1. 1.
    Initial Embedding Normalization:
    $y_{o}=\mathrm{Norm}(x_{o}),$
  2. 2.
    Input- & Output-Normalization per Layer:
    $y_{l}=x_{l}+\mathrm{Norm}\Bigl{(}\mathrm{Module}\bigl{(}\mathrm{Norm}(x_{l})\bigr{)}\Bigr{)},$
  3. 3.
    Final Embedding Normalization:
    $y_{L}=\mathrm{Norm}(x_{L}),$
where $x_{o}$ denotes the output of the embedding layer, the hidden input state. $y_{0}$ represents the normalized input hidden state. $x_{L}$ denotes the hidden state output by the final layer $L$ of the Transformer sub-layer. This design unifies pre- and output-normalization to regulate variance from both ends. For clarity, the locations of normalization layers in the Post-, Pre-, and Peri-LN architectures are illustrated in Figure 2.
Controlling Variance & Preserving Gradients.
By normalizing both the input and output of each sub-layer, Peri-LN constrains the residual spikes common in Pre-LN, while retaining a stronger gradient pathway than Post-LN. Concretely, if $\mathrm{Norm}(\mathrm{Module}(\mathrm{Norm}(x_{l})))$ has near-constant variance $\beta_{0}$, then
$\mathrm{Var}(x_{l+1})\;\approx\;\mathrm{Var}(x_{l})+\beta_{0},$
leading to linear or sub-exponential hidden state growth rather than exponential blow-up. We empirically verify this effect in Section 4.4.2.
(a) Learning rate exploration
(b) Training loss
(c) Gradient-norm
Figure 3: Performance comparison of Post-LN, Pre-LN, and Peri-LN Transformers during pre-training. Figure 3(a) llustrates the pre-training loss across learning rates. Pre-training loss and gradient norm of best performing $400$M size Transformers are in Figure 3(b) and 3(c). Consistent trends were observed across models of different sizes.
(a) Training loss
(b) Loss in the final $5$B token interval
(c) Gradient-norm
Figure 4: Freezing learnable parameter $\gamma$ of output normalization layer in Peri-LN. we set $\gamma$ to its initial value of $1$ and keep it fixed.
Open-Sourced Peri-LN Models: Gemma2222 & OLMo2222.
Both Gemma2222 and OLMo2222, which apply output layer normalization, employ the same peri-normalization strategy within each Transformer layer. However, neither model rigorously examines how this placement constrains variance or mitigates large residual activations. Our work extends Gemma2222 and OLMo2222 by offering both theoretical and empirical perspectives within the Peri-LN scheme. Further discussion of the OLMo2222 is provided in Appendix G.
3.4 Stability Analysis in Normalization Strategies
We analyze training stability in terms of the magnitude of activation. To this end, we examine the gradient norm with respect to the weight of the final layer in the presence of massive activation. For the formal statements and detailed proofs, refer to Appendix D.
Proposition 3.1 (Informal).
Let $\mathcal{L}(\cdot)$ be the loss function, and let $W^{(2)}$ denote the weight of the last layer of $\mathrm{MLP}(\cdot)$. Let $\gamma$ be the scaling parameter in $\mathrm{Norm}(\cdot)$, and let $D$ be the dimension. Then, the gradient norm for each normalization strategy behaves as follows.
(1) Pre-LN (exploding gradient). Consider the following sequence of operations:
$\tilde{x}=\mathrm{Norm}(x),a=\mathrm{MLP}(\tilde{x}),o=x+a,$ (3)
then
$\left\lVert\frac{\partial\mathcal{L}(o)}{\partial W_{i,j}^{(2)}}\right\rVert\;\propto\;\|h_{i}\|,$ (4)
where $h:=\mathrm{ReLU}\left(\tilde{x}W^{(1)}+b^{(1)}\right)$. In this case, when a massive activation $\|h\|$ occurs, an exploding gradient $\|\partial\mathcal{L}/\partial W^{(2)}\|$ can arise, leading to training instability.
(2) Peri-LN (self-regularizing gradient). Consider the following sequence of operations:
$\tilde{x}=\mathrm{Norm}(x),a=\mathrm{MLP}(\tilde{x}),\tilde{a}=\mathrm{Norm}(a),o=x+\tilde{a},$ (5)
then
$\left\lVert\frac{\partial\mathcal{L}(o)}{\partial W_{i,j}^{(2)}}\right\rVert\;\leq\;\frac{4\,\gamma\,\sqrt{D}\,\|h\|}{\|a\|},$ (6)
where $h:=\mathrm{ReLU}\left(\tilde{x}W^{(1)}+b^{(1)}\right)$. In this case, even when a massive activation $\|h\|$ occurs, $\mathrm{Norm}(\cdot)$ introduces a damping factor $\|a\|$, which ensures that the gradient norm $\|\partial\mathcal{L}/\partial W^{(2)}\|$ remains bounded.
(3) Post-LN (vanishing gradient). Consider the following sequence of operations:
$a=\mathrm{MLP}(x),o=x+a,\tilde{o}=\mathrm{Norm}(o),$ (7)
then
$\left\lVert\frac{\partial\mathcal{L}(\tilde{o})}{\partial W_{i,j}^{(2)}}\right\rVert\;\leq\;\frac{4\,\gamma\,\sqrt{D}\,\|h\|}{\|x+a\|},$ (8)
where $h:=\mathrm{ReLU}\left(xW^{(1)}+b^{(1)}\right)$. In this case, when a massive activation $\|h\|$ occurs, $\mathrm{Norm}(\cdot)$ introduces an overly suppressing factor $\|x+a\|$, which contains a separate huge residual signal $x$, potentially leading to a vanishing gradient $\|\partial\mathcal{L}/\partial W^{(2)}\|$.
We have compiled a Table 1 that provides a overview of the variance and gradient intuition for each layer normalization strategy.
Table 2: Average benchmark scores (with standard deviations) across $5$ different training seeds for Post-, Pre-, and Peri-Layer Normalization language models. Each model size excludes the embedding parameters. Loss denotes the evaluation loss on random samples of the C4444 dataset (Raffel et al., 2020). Arch. denotes architecture, and Avg. denotes the averaged benchmark score across tasks. SFT avg. denotes the averaged benchmark score across tasks of instruction fine-tuned models. When calculating the evaluation score, diverged checkpoints were excluded.
Size Arch. ARC-Easy HellaSwag PIQA SIQA Winogrande Avg. $\uparrow$ Loss $\downarrow$ SFT Avg. $\uparrow$
Post-LN $35.70$$\pm 1.09$ $28.91$$\pm 0.16$ $62.26$$\pm 0.73$ $34.48$$\pm 1.04$ $50.88$$\pm 0.75$ $42.45$ $7.46$ $46.44$
$400$M Pre-LN $54.87$$\pm 1.63$ $34.17$$\pm 1.66$ $68.79$$\pm 1.34$ $39.73$$\pm 0.59$ $50.88$$\pm 2.35$ $49.69$ $3.43$ $49.96$
Peri-LN 57.51 ±$\pm 0.81$ 37.46 ±$\pm 0.34$ 69.48 ±$\pm 0.39$ 40.64 ±$\pm 0.51$ 52.74 ±$\pm 0.67$ 51.57 3.34 51.96
Post-LN $42.92$$\pm 0.93$ $31.69$$\pm 0.41$ $66.72$$\pm 0.40$ $35.84$$\pm 0.61$ $50.30$$\pm 1.87$ $45.49$ $5.38$ $48.95$
$1.5$B Pre-LN $61.51$$\pm 1.22$ $39.88$$\pm 1.53$ $71.41$$\pm 0.88$ $41.23$$\pm 0.97$ $54.51$$\pm 2.07$ $53.71$ $3.29$ $53.89$
Peri-LN 66.17 ±$\pm 0.21$ 43.94 ±$\pm 0.34$ 73.63 ±$\pm 0.24$ 42.34 ±$\pm 0.83$ 56.64 ±$\pm 0.44$ 56.55 3.18 56.94
Post-LN $45.30$$\pm 3.23$ $33.59$$\pm 0.44$ $66.45$$\pm 2.86$ $35.82$$\pm 1.09$ $51.10$$\pm 1.60$ $46.45$ $4.43$ $49.33$
$3.2$B Pre-LN $65.24$$\pm 2.32$ $44.23$$\pm 2.32$ $73.86$$\pm 1.19$ $42.68$$\pm 0.07$ $57.42$$\pm 2.51$ $56.69$ $3.20$ $57.08$
Peri-LN 68.73 ±$\pm 0.57$ 46.99 ±$\pm 0.21$ 74.31 ±$\pm 0.41$ 43.00 ±$\pm 0.73$ 59.76 ±$\pm 0.78$ 58.56 3.11 59.02
4 Experiments
In this section, we provide a comprehensive empirical comparison of Post-, Pre-, and Peri-Layer Normalization (LN) across large-scale Transformer pre-training and subsequent evaluations on the language domain.
4.1 Experimental Setting
Excluding the embedding parameters, model sizes were set to $400$M, $1.5$B, and $3.2$B parameters, respectively. Each model was trained on $30$ billion tokens. For a fair comparison, we pre-trained each model with five different training seeds. We conducted a systematic exploration of learning rates, ranging from $1\times 10^{-4}$ to $5\times 10^{-3}$. The global batch size is set to $256$. We use Adam optimizer with cosine learning rate scheduler. The sequence length was set to $8192$, and the weight decay coefficient was fixed at $0.033$. We employed Megatron-LM [^4] https://github.com/NVIDIA/Megatron-LM to pre-train the Transformers under each layer normalization strategy. We used the DCLM-baseline dataset (Li et al., 2024a) for our experiments, along with the “cl100100100100k_base” version of the TikToken tokenizer [^5] https://github.com/openai/tiktoken. Unless otherwise noted, most training and model settings followed those of the DCLM experiments. For layer normalization, we primarily employed RMSNorm. Further details are provided in Appendix B.
4.2 Pre-Training Large Language Models
Figure 3(a) illustrates the pre-training loss across learning rates for models ranging in size from $400$M to $3.2$B parameters. Notably, the Peri-LN architecture consistently achieves superior loss curves over this entire model size. Since Pre-LN shows best performance at learning rate $2\times 10^{-3}$ across all model size, we set this to the default learning rate for Pre-LN and Peri-LN. Unlike Pre-LN, Post-LN’s appropriate learning rate lies in a lower range, so we have provided a separate summary in Appendix C.1. In Figures 3(b) and 3(c), we compare the pre-training loss and the gradient norm curve at each LN strategy’s best-performing learning rate of $400$M size models. The same trend was observed across different model sizes. In particular, when sweeping over training seeds, Pre-LN and Post-LN exhibited many spikes in the gradient norm curve, whereas Peri-LN showed relatively few spikes, supporting Proposition 3.1. This reduction is consistent across both large and small learning rates.
We provide additional results in Appendix F, including experiments using LayerNorm (instead of RMSNorm), analyses of varying sequence lengths, the effect of different training-token counts, and an ablation study on embedding LN. We also investigate pre-training using the OLMo2222-style Peri-LN architecture in Appendix G. These findings are likewise consistent with those previously reported.
(a) Absolute magnitude growth
(b) Variance growth
Figure 5: This figure shows the forward growth patterns of hidden states for different architectures, highlighting the structural impact of normalization placement. Each model contains $1.5$ billion parameters (excluding the embedding size). We confirmed that the observed trend remains consistent across all model sizes.
(a) Grad-norm at init
(b) Grad-norm at final
(c) Grad-variance at init
(d) Grad-variance at final
Figure 6: Comparison of the backward gradient norm and variance for Post-LN, Pre-LN, and Peri-LN Transformers at initialization and at the final stage. Model size is $1.5$B. We confirmed that the observed trend remained consistent across all model sizes. init denotes the initialization of the model parameters.
4.3 Benchmark Evaluations & Instruction Tuning
To evaluate how well the pre-trained model’s training loss aligns with its benchmark performance, we conducted five separate benchmark evaluations. Furthermore, to examine how supervised fine-tuning (SFT) improves scores and instruction-following capabilities under different layer normalization strategies in Transformer architectures, we performed additional training using the LIMA dataset (Ouyang et al., 2022; Zhou et al., 2023). Diverged checkpoints were excluded when calculating the evaluation score (mostly occurs in Pre-LN). Additional training hyperparameters for SFT are given in Appendix B.2. As shown in Table 2, Peri-LN consistently demonstrates superior performance across all model sizes. Additionally, we note that, beyond the improved scores, the standard deviation of the benchmark results across different training seeds is reduced by more than half with Peri-LN. From this, we observe that Peri-LN helps maintain consistency not only in gradient stability and final loss but also in benchmark performance. For the evaluation loss, we used $10$K random samples from the C4444 dataset (Raffel et al., 2020). Detailed configurations and individual scores are provided in Appendix H.
4.4 Systematic Analysis
Despite emerging empirical evidence that Peri-LN can stabilize training and bolster performance, many open questions remain: Which design factors are crucial for reaping its benefits at scale (§4.4.1)? How does it regulate hidden-state variance and gradient flow more effectively (§4.4.2, §4.4.3)? And why might certain variants outperform standard LN placements in large Transformer architectures (§4.4.4)? Addressing these points, we present a systematic analysis of Peri-LN’s mechanics at following sub-sections.
4.4.1 Learnable Parameter $\gamma$ of RMSNorm
To investigate the impact of module output normalization on training stability, as proposed in the Proposition 3.1, we fixed the learnable parameter $\gamma$ of RMSNorm to $1$, isolating the effect of normalization. As shown in Figure 4, we observe that simply normalizing the output of each module reduces gradient spikes and provides improvements in the loss comparing to the pre-normalization only transformers. Nonetheless, we also confirmed that allowing $\gamma$ to be learnable yields slightly better performance. We confirm that this trend remains consistent in both the $400$M and $1.5$B models. In this experiment, we omitted Peri-LN’s embedding layer normalization in order to isolate and evaluate the precise role and benefits of output layer normalization.
(a) At initialization
(b) Learnable scale $\gamma$ in Output-LN
(c) After $30$B tokens training
(d) Frozen $\gamma$ in Output-LN
Figure 7: Angular distance of hidden state is presented in Figure  7(a)7(c), and 7(d). In Figure 7(b), we monitor $\gamma$ of every Output-LN in Peri-LN during training. We use $30$B tokens trained $400$M size model in this experiments.
4.4.2 Growth of Hidden State
To examine in greater depth how Peri-LN affects the forward propagation, we analyzed the absolute magnitude and variance of the hidden states using $1,000$ samples from the Wikitext dataset (Merity et al., 2016). Figure 5 illustrates how normalization strategies influence the forward-path hidden states over the course of training.
Two particular aspects of hidden-state behavior were noted: First, with respect to model depth, Post-LN maintains stable hidden-state magnitude and variance owing to the presence of a normalization layer in the main path (“Highway”). By contrast, because Pre-LN does not normalize the outputs of each Attention and MLP module, the magnitude and variance of the hidden states exhibit exponential growth following the addition operation in the residual path. For the Peri-LN architecture, which leverages Output-LN, we observed that hidden states remain comparatively well-managed. Next, regarding changes over training iterations, Post-LN continues to normalize each block’s output, which prevents any substantial increase or decrease in the trend as training progresses. Meanwhile, Pre-LN exhibits a relatively linear variance distribution at initialization, but escalates exponentially to extremely large values over successive iterations. In contrast, Peri-LN fluctuates more moderately, owing to the role of Output-LN in controlling hidden-state magnitude and variance throughout training.
4.4.3 Layer-wise Gradient Norm & Variance
Ensuring a uniform gradient flow in large-scale model training is crucial for balanced learning across the entire network (Yang & Hu, 2021; Yang et al., 2024). As shown in Figure 6, in Post-LN, gradients decrease as they propagate backward through the layers in the final stage of training, which can lead to vanishing gradients in lower-index layers. In Pre-LN, gradients increase as they propagate backward through the layers at initialization, potentially causing explosive gradients in the early phase of training. Both strategies display non-uniform gradient distributions—either vanishing or exploding—at different stages of training. On the other hand, Peri-LN demonstrates a consistent, layer-wise gradient distribution at both initialization and the end of training. By maintaining comparatively uniform gradients with lower variance across layers, Peri-LN avoids the extremes of vanishing or exploding behaviors. This stability is particularly beneficial in deeper architectures, where balanced gradient flow is essential for effective backpropagation.
4.4.4 Hidden State Representation
We utilize angular distance (Li et al., 2024b) as a metric to assess redundancy between hidden states at initialization and after training. This approach allows us to quantify how similar or distinct the representations were across layers. As shown in Figure 7(a), Post-LN exhibits smaller angular distances due to the LN being located on the main path, whereas Pre-LN and Peri-LN begin with very similar states. As shown in Figure 7(c), at the end of training, Pre-LN tends to produce more redundant hidden state representations compared to the others. This phenomenon may stem from Pre-LN’s repeated residual additions, which amplify certain representations over others.
To investigate this further, we focus on module output normalization, which is the main distinguishing factor between Pre-LN and Peri-LN. As shown in Figure 7(b), the learnable scale starts around $1$ in the early stages of training and gradually changes with increasing depth. Because Peri-LN preserves the identity path, it appears to adjust the scale of the module output accordingly. This suggests that the exponential growth of the main path’s magnitude in Pre-LN diminishes the relative contribution of individual modules, resulting in more redundant hidden representations. To support this, Figure 7(d) shows that fixing the learnable scale of Peri-LN’s module output LN at $1$ causes the main path contribution to decrease in deeper layers. This finding confirms the role of module output normalization in controlling hidden state redundancy.
5 Conclusion
We explore the placement of layer normalization within the Transformer architecture to better understand its role during training. By systematically comparing Post-LN, Pre-LN, and Peri-LN, we highlight their distinct impacts on stability, gradient propagation, and optimization dynamics. Our findings suggest that placing LN on module outputs in addition to the Pre-LN can help manage large activations while preserving beneficial gradient flow, thereby offering a promising balance for stable optimization. By unifying these approaches under the term Peri-LN, we seek to consolidate existing variants and encourage deeper investigation into this underexplored alternative.
Acknowledgements
We thank our colleague Jeongin Bae for inspiring the underlying motivation for this research. We are also grateful to Jung Hyun Lee, Seonghyeon Kim, and Seunghyun Seo for their valuable assistance during the early stages of our research discussions. Finally, we extend our gratitude to Gichang Lee, Lead of the Backbone Mission at NAVER Cloud, for his unwavering support in connecting us with key resources throughout this research.
Impact Statement
The rapid advancement of Transformer-based large language models (LLMs) has enabled remarkable breakthroughs in natural language understanding and generation. However, these models also pose significant challenges, including concerns around safety, bias, and the computational cost associated with large-scale training. As LLMs become increasingly integral to various AI applications, ensuring their stability, efficiency, and accessibility remains a critical research focus.
Our work addresses these challenges by proposing a more stable and cost-effective large-scale training methodology. By improving training efficiency and reducing the associated computational overhead, we lower the barrier to entry for organizations seeking to develop or fine-tune foundation models. This democratization of LLM technology fosters broader participation in AI research and development, accelerating innovation while mitigating concerns over resource concentration in a few major players. Given the growing industry focus on optimizing LLM deployment costs, our contributions are particularly relevant in the current AI research landscape.
Another positive externality of improving the cost-effectiveness of large-scale training is the reduction in the environmental impact of AI research. Training state-of-the-art LLMs requires vast computational resources, leading to significant energy consumption and carbon emissions. By enhancing the efficiency of large-scale training, our approach helps mitigate these environmental costs, making AI research more sustainable. Lowering the energy footprint of training not only aligns with global efforts toward greener computing but also enables a wider range of institutions, including smaller research labs and academic groups, to participate in cutting-edge AI development without excessive resource constraints. This shift promotes a more inclusive and responsible AI ecosystem, ensuring that advancements in foundation models are both accessible and environmentally conscious.
Appendix A Related Work
Activation Dynamics in Large Language Models.
Studies on the distribution and magnitude of activations in deep neural networks have revealed that certain outlier features can significantly affect model behavior and efficiency. Dettmers et al. (2022) examined Transformer architectures, highlighting how specific feature dimensions may exhibit unusually large values (outliers) that disrupt quantization and overall system performance. Extending this line of work, Sun et al. (2024) identified the occurrence of “massive activations”—extremely large activation values that persist across multiple layers. Unlike standard outliers, these massive activations remain relatively invariant to different inputs, effectively functioning as implicit bias terms in large language models (LLMs). Notably, such extreme values can skew the self-attention mechanism, causing the model to attend disproportionately to certain tokens. These observations demonstrate that even with standard normalization layers in place, hidden biases may linger in internal representations, underscoring the importance of deeper analyses of activation behavior in LLMs.
Variance Control and Normalization in Convolutional Networks.
The interplay between activation variance and training stability has also been extensively explored in convolutional neural networks (CNNs). De & Smith (2020) showed that Batch Normalization (BN) stabilizes the training of residual networks by effectively downscaling activation variance in the residual branches, thereby improving gradient behavior. However, BN imposes certain constraints, such as dependence on batch size and additional computational overhead for estimating batch statistics. Consequently, several normalization-free or alternative normalization approaches have been investigated. For instance, Brock et al. (2021b) introduced “Normalizer-Free ResNets,” which manage activation variance through learnable scaling parameters. This approach achieved competitive performance without relying on BN, highlighting the critical role of effective variance control in fostering stable optimization and strong generalization in CNNs.
Layer Normalization in Transformers.
Training stability in Transformer architectures is closely tied to the choice and placement of layer normalization (LN). Xiong et al. (2020) reported that Transformers employing a Post-LN structure often suffer from gradient instabilities at initialization, requiring a careful learning-rate warm-up phase to mitigate these issues. In contrast, Pre-LN helps maintain more stable gradients during the early stages of training. However, Kedia et al. (2024) showed that while Post-LN can lead to vanishing or exploding gradients in deep Transformers, Pre-LN may induce hyperbolic gradient growth. These findings illustrate the nuanced trade-offs of normalization placement and draw parallels to earlier CNN studies, where careful management of activation variance proved essential for stable deep learning.
Gradient Propagation and Depth Scaling
Ensuring consistent gradient propagation across many layers is pivotal for stable training in very deep models. Yang & Hu (2021) (Tensor Programs IV) introduced the concept of Maximal Update Parametrization ($\mu$P) in the infinite-width regime to preserve feature learning, preventing gradients from collapsing into kernel-like dynamics. Building on this, Yang et al. (2024) (Tensor Programs VI) proposed Depth-$\mu$P, which scales residual branches and learning rates according to network depth. Their theoretical analysis indicates that improper depth-dependent scaling leads to vanishing or exploding gradients, ultimately diminishing the diversity of learned representations. These insights highlight the necessity for principled scaling strategies and careful initialization to maintain robust gradient flow in deep architectures.
Summary.
Taken together, these studies underscore the importance of managing activation variance and hidden biases to achieve stable training and expressive internal representations in modern deep networks. In Transformer-based models, normalization choice and placement—such as Post-LN, Pre-LN, or other variants—play a significant role in controlling gradient dynamics and overall performance. While Post-LN and Pre-LN have received significant attention, we focus on a comparative analysis that includes Peri-LN, an alternative normalization placement that has thus far been underexplored but holds potential for enhancing training stability and model performance.
Appendix B Detailed Experimental Setting
In this section, we provide detailed configurations of both the pretraining and supervised fine-tuning to reproduce our results.
B.1 Configurations on Pre-Training
The common training settings are provided in Table 3. Embedding settings for the language models are listed in Table 4. For the model architecture, we primarily follow the Llama $3$ architecture (Dubey et al., 2024). In the MLP module, we use SwiGLU activations. Additional details regarding the model configurations are shown in Table 5. Note that embedding parameters are excluded from the model size. Unless otherwise noted, most training and model settings follow those of the DCLM experiments (Li et al., 2024a). During the pretraining stage, each model was trained under a controlled random seed.
Table 3: Common configurations. LR Schedule denotes learning rate schedule.
Global Batch Size Weight Decay Iterations Optimizer LR Schedule Warmup
$256$ $0.033$ $14400$ Adam Cosine $10$%
Table 4: Embedding configurations.
Max Position Embeddings Position Embedding Type Untie-embeddings-and-output-weights
$8192$ Rope $True$
Table 5: Model configurations.
Size $n_{layers}$ $n_{heads}$ $d_{model}$ $d_{head}$
$400$M $24$ $8$ $1024$ $128$
$1.5$B $24$ $16$ $2048$ $128$
$3.2$B $32$ $16$ $2560$ $160$
B.2 Configurations on Supervised Fine-Tuning
To examine downstream task performance after instruction tuning, we employed a high-quality LIMA alignment training set consisting of $1,000$ samples (Zhou et al., 2023). Our supervised fine-tuning configuration was slightly modified from the original setup of LIMA: we fine-tuned the model for $15$ epochs with a batch size of $128$. The optimizer was Adam with an initial learning rate of $1$e-$5$ and a cosine learning rate schedule. We selected the best checkpoints for each model by evaluating on OpenBookQA (Mihaylov et al., 2018), CommonSenseQA (Talmor et al., 2019), and the NLI dataset in GLUE (Wang et al., 2018).
Appendix C Additional Results on Pre-Training Study
C.1 Post-Layer Normalization Architecture & Learning Rate Exploration
In order to identify the optimal performance configuration for Post-LN within the experimental setup, we conducted a learning rate exploration as shown in Figure 8. Because the appropriate learning rate for Post-LN fell into a much lower range than those for Pre-LN and Peri-LN, we treated it separately. For each Post-LN setting, the best learning rate was determined as the one yielding the lowest final training loss, with the random seed held constant during this selection process.
Figure 8: Learning rate explorations for Post-LN architectures.
C.2 Best Performing Checkpoints Comparisons of Other Model Sizes
As an extension of Section 4.2, we present below the results for additional model sizes that were omitted previously due to space constraints. In Figures 9, we compare the pre-training loss and the gradient norm curve at each LN strategy’s best-performing learning rate of $3.2$B and $1.5$B size models.
(a) Training loss for $3.2$B
(b) Gradient-norm for $3.2$B
(c) Training loss for $1.5$B
(d) Gradient-norm for $1.5$B
Figure 9: Performance comparison of Post-LN, Pre-LN, and Peri-LN Transformers during pre-training for other two.
C.3 Gradient Spikes during Pre-Training
Although this observation is empirical, it remains sufficiently noteworthy to discuss. During extensive pre-training ablation experiments, we frequently observed gradient-norm spikes. Specifically, when examining how these spikes vary with different learning rates, we found that Pre-LN exhibited irregular spikes regardless of the learning rate setting, with a particularly high occurrence in the early stages of training. In contrast, such spikes were rarely observed under Peri-LN.
Appendix D Proof of Theoretical Insight
To support the claim that Peri-LN enhances the stability of training in such cases, we analyze the gradient norm in the final layer. For simplicity, we use RMSNorm and ReLU here.
Proposition D.1.
Consider the following sequence of operations:
$\quad\tilde{x}=\mathrm{RMSNorm}(x),$
$\quad a=\mathrm{ReLU}(\tilde{x}W^{(1)}+b^{(1)})W^{(2)}+b^{(2)},$
$\quad o=x+a.$
Then,
$\frac{\partial\mathcal{L}(o)}{\partial W_{i,j}^{(2)}}=h_{i}(\hat{p}_{j}-y_{j}),$ (9)
where $h:=\mathrm{ReLU}\left(xW^{(1)}+b^{(1)}\right)$, $\hat{p}:=\mathrm{softmax}(o)$, and $y$ is the label (one-hot vector).
Proof.
By the chain rule,
$\frac{\partial\mathcal{L}(o)}{\partial W_{i,j}^{(2)}}=\underbrace{\frac{\partial\mathcal{L}(o)}{\partial o}}_{\text{(a)}:1\times D}\times\underbrace{\frac{\partial o}{\partial a}}_{\text{(b)}:D\times D}\times\underbrace{\frac{\partial a}{\partial W_{i,j}^{(2)}}}_{\text{(c)}:D\times 1}.$ (10)
(a) It is known that
$\frac{\partial\mathcal{L}(o)}{\partial o_{k}}=\hat{p}_{k}-y_{k}.$ (11)
So,
$\frac{\partial\mathcal{L}(o)}{\partial o}=\begin{bmatrix}\hat{p}_{1}-y_{1}&\hat{p}_{2}-y_{2}&\cdots&\hat{p}_{D}-y_{D}\end{bmatrix}.$ (12)
(b) Since $o=x+a$,
$\frac{\partial o}{\partial a}=I.$ (13)
(c) Recall that
$a:=\mathrm{ReLU}(\tilde{x}W^{(1)}+b^{(1)})W^{(2)}+b^{(2)}.$ (14)
For convenience, let
$h:=\mathrm{ReLU}(\tilde{x}W^{(1)}+b^{(1)}).$ (15)
Then, we have
$\frac{\partial a_{k}}{\partial W_{i,j}^{(2)}}=\frac{\partial}{\partial W_{i,j}^{(2)}}\left(\sum_{p=1}^{H}h_{p}W_{p,k}^{(2)}+b_{k}^{(2)}\right)=h_{i}\,\delta_{k,j}.$ (16)
In vector representation,
$\frac{\partial a}{\partial W_{i,j}^{(2)}}=\begin{bmatrix}0&\cdots&h_{i}&\cdots&0\end{bmatrix}^{\top},$ (17)
where the only nonzero entry is in the $j$-th component.
Thus, by putting these all together,
$\frac{\partial\mathcal{L}(o)}{\partial W_{i,j}^{(2)}}=h_{i}(\hat{p}_{j}-y_{j}).$ (18)
Proposition D.2.
Consider the following sequence of operations:
$\quad\tilde{x}=\mathrm{RMSNorm}(x),$
$\quad a=\mathrm{ReLU}(\tilde{x}W^{(1)}+b^{(1)})W^{(2)}+b^{(2)},$
$\quad\tilde{a}=\mathrm{RMSNorm}(a),$
$\quad o=x+\tilde{a}.$
Then,
$\left\lVert\frac{\partial\mathcal{L}(o)}{\partial W_{i,j}^{(2)}}\right\rVert\leq\frac{4\gamma\sqrt{D}\|h\|}{\|a\|},$ (19)
where $\gamma$ is the scaling parameter used in $\mathrm{RMSNorm}(\cdot)$, $D$ is the dimensionality, and $h:=\mathrm{ReLU}\left(xW^{(1)}+b^{(1)}\right)$.
Proof.
By the chain rule,
$\frac{\partial\mathcal{L}(o)}{\partial W_{i,j}^{(2)}}=\underbrace{\frac{\partial\mathcal{L}(o)}{\partial o}}_{\text{(a)}:1\times D}\times\underbrace{\frac{\partial o}{\partial\tilde{a}}}_{\text{(b)}:D\times D}\times\underbrace{\frac{\partial\tilde{a}}{\partial a}}_{\text{(c)}:D\times D}\times\underbrace{\frac{\partial a}{\partial W_{i,j}^{(2)}}}_{\text{(d)}:D\times 1}.$ (20)
(a) We have
$\left\|\frac{\partial\mathcal{L}(o)}{\partial o}\right\|=\|\hat{p}-y\|\leq\|\hat{p}\|+\|y\|=2.$ (21)
(b) We also have
$\left\|\frac{\partial o}{\partial\tilde{a}}\right\|=\|I\|=1.$ (22)
(c) Recall that
$\tilde{a}:=\mathrm{RMSNorm}(a)=\gamma\cdot\frac{a}{\sqrt{\frac{1}{D}\sum_{k=1}^{D}a_{k}^{2}+\epsilon}}.$ (23)
Then, $\frac{\partial\tilde{a}}{\partial a}$ is the Jacobian matrix $J$ of $\mathrm{RMSNorm}(\cdot)$. For brevity, let
$\alpha:=\frac{1}{D}\sum_{k=1}^{D}(a_{k})^{2}.v$ (24)
Then,
$J_{p,q}=\frac{\partial\tilde{a}_{p}}{\partial a_{q}}$ $=\gamma\cdot\frac{\partial}{\partial a_{q}}\left(\frac{a_{p}}{\sqrt{\alpha+\epsilon}}\right)$ (25)
$=\gamma\cdot\frac{1}{\sqrt{\alpha+\epsilon}}\frac{\partial a_{p}}{\partial a_{q}}+\gamma\cdot a_{p}\frac{\partial}{\partial a_{q}}\left(\frac{1}{\sqrt{\alpha+\epsilon}}\right)$ (26)
$=\gamma\cdot\frac{1}{\sqrt{\alpha+\epsilon}}\delta_{p,q}-\gamma\cdot\frac{a_{p}a_{q}}{D(\alpha+\epsilon)^{3/2}}.$ (27)
In matrix representation,
$J=\underbrace{\frac{\gamma}{\sqrt{\alpha+\varepsilon}}I}_{A}-\underbrace{\frac{\gamma}{D(\alpha+\varepsilon)^{3/2}}\left(a\right)^{\top}\left(a\right)}_{B}.$ (28)
Then, we have
$\|A\|=\left\|\frac{\gamma}{\sqrt{\alpha+\varepsilon}}I\right\|=\frac{\gamma}{\sqrt{\alpha+\varepsilon}}\|I\|=\frac{\gamma}{\sqrt{\alpha+\varepsilon}},$ (29)
and
$\|B\|=\left\|\frac{\gamma}{D(\alpha+\varepsilon)^{3/2}}\left(a\right)^{\top}\left(a\right)\right\|=\frac{\gamma}{D(\alpha+\varepsilon)^{3/2}}\times D\alpha=\frac{\gamma\alpha}{(\alpha+\varepsilon)^{3/2}}.$ (30)
So, we have
$\|J\|=\|A-B\|\leq\|A\|+\|B\|\leq\frac{2\gamma}{\sqrt{\alpha}}=\frac{2\gamma\sqrt{D}}{\|a\|}.$ (31)
(d) Since
$\frac{\partial a}{\partial W_{i,j}^{(2)}}=\begin{bmatrix}0&\cdots&h_{i}&\cdots&0\end{bmatrix}^{\top},$ (32)
we have
$\left\|\frac{\partial a}{\partial W_{i,j}^{(2)}}\right\|\leq\|h\|.$ (33)
Thus,
$\left\|\frac{\partial\mathcal{L}(o)}{\partial W_{i,j}^{(2)}}\right\|\leq 2\times 1\times\frac{2\gamma\sqrt{D}}{\|a\|}\times\|h\|=\frac{4\gamma\sqrt{D}\|h\|}{\|a\|}.$ (34)
Proposition D.3.
Consider the following sequence of operations:
$\quad a=\mathrm{ReLU}(xW^{(1)}+b^{(1)})W^{(2)}+b^{(2)},$
$\quad o=x+a,$
$\quad\tilde{o}=\mathrm{RMSNorm}(o).$
Then,
$\left\lVert\frac{\partial\mathcal{L}(\tilde{o})}{\partial W_{i,j}^{(2)}}\right\rVert\leq\frac{4\gamma\sqrt{D}\|h\|}{\|x+a\|},$ (35)
where $\gamma$ is the scaling parameter used in $\mathrm{RMSNorm}(\cdot)$, $D$ is the dimensionality, and $h:=\mathrm{ReLU}\left(xW^{(1)}+b^{(1)}\right)$.
Proof.
The proof is analogous to the proof of the previous proposition. ∎
Appendix E Additional Results on Growth of Hidden State
In this section, we examine the $400$M- and $3.2$B-parameter models, which were omitted in Section 4.4.2 due to space constraints. As illustrated in Figures 10 and 11, these models exhibit the same overall trend.
(a) Absolute magnitude growth
(b) Variance growth
Figure 10: The forward growth patterns of hidden state for different architectures highlight the structural impact of normalization placement. $3.2$B size model.
(a) Absolute magnitude growth
(b) Variance growth
Figure 11: The forward growth patterns of hidden state for different architectures highlight the structural impact of normalization placement. $400$M size model.
Appendix F Additional Experimental Results on Ablation Study
F.1 Amount of Training Tokens
In order to investigate whether the learning behavior of each LN strategy varies with the number of training tokens, we conducted an additional round of learning-rate exploration for both the Pre-LN and Peri-LN architectures. As shown in Figure 12, even as the number of training tokens increases, there is no observable shift in the optimal learning-rate range. Based on these findings, we conclude that our overall results remain consistent, even when the training token count is further increased. Furthermore, although a learning rate of $5\times 10^{-3}$ leads to divergence in the smaller-scale experiments with $8$B or $16$B training tokens, it does not do so in the $30$B-token setting. We attribute this discrepancy to the $10$% warmup rate, suggesting that the warmup phase may be insufficient for the smaller-scale experiments.
Figure 12: Learning rate explorations of Pre-& Peri-LN architecture with sequence length $2048$ configuration.
F.2 Sequence Length
In language models, the number of iterations per token is influenced by the sequence length, which in turn, along with the batch size, affects training statistics. We conducted an experiment to determine whether the performance trend changes when the sequence length is reduced from 8192, as set in the main text, to 2048. As shown in Figure 13, Peri-LN still surpasses Pre-LN in the learning rate exploration.
Figure 13: Learning rate explorations of Pre-& Peri-LN architecture with sequence length $2048$ configuration.
F.3 Warm-up
Warmup is widely recognized to influence training stability. To investigate whether a $10$% warmup rate might unfairly disadvantage Pre-LN, we conducted an additional learning-rate exploration using a $30$% warmup rate. As illustrated in Figure 14, the overall trend remained unchanged, and Peri-LN continued to exhibit better performance than Pre-LN in terms of loss. Furthermore, we observed that increasing the warmup rate from $10$% to $30$% did not reduce the frequency of gradient norm spikes in Pre-LN.
Figure 14: Learning rate explorations of Pre-& Peri-LN architecture with warmup $30$% configuration.
F.4 RMSNorm & LayerNorm
As illustrated in Figure 15, we conducted experiments in which RMSNorm and LayerNorm were interchanged. Consistent with the findings reported in (OLMo et al., 2024), we did not observe any notable performance differences in our RMSNorm and LayerNorm replacement experiments. Learning rate was set to $2$e-$3$ (best performance learning rate).
(a) Pre-training loss curve
(b) Gradient-norm curve
Figure 15: LayerNorm vs. RMSNorm on Peri-LN architecture. $400$M size model.
F.5 Embedding Layer Normalization of Peri-Layer Normalization Transformers
Motivated by Takase et al. (2023), we empirically explore the addition of embedding layer normalization to improve training stability and overall model performance in Transformer architectures. As illustrated in Figures 16, 17, and 18, incorporating Embedding LN in the Peri-LN architecture yields a slight improvement in pre-training loss. Furthermore, our empirical observations suggest that this effect becomes more pronounced in smaller models.
(a) Pre-training loss curve
(b) Gradient-norm curve
Figure 16: Loss and Gradient-norm curves comparing the presence and absence of Embedding LN in the Peri-LN architecture. $400$M size model.
(a) Pre-training loss curve
(b) Gradient-norm curve
Figure 17: Loss and Gradient-norm curves comparing the presence and absence of Embedding LN in the Peri-LN architecture. $1.5$B size model.
(a) Pre-training loss curve
(b) Gradient-norm curve
Figure 18: Loss and Gradient-norm curves comparing the presence and absence of Embedding LN in the Peri-LN architecture. $3.2$B size model.
Appendix G Output-Layer Normalization with QK-Norm Architecture
Query and Key layer-normalization (QK-Norm) has been widely studied in modern Transformer architectures (Wortsman et al., 2024; Zhai et al., 2023; OLMo et al., 2024). In particular, OLMo et al. (2024) reported that QK-Norm combined with module output layer-normalization (output-LN, $B$ in Figure 19 referred to as “reordered norm” in the OLMo2222 paper) improves both training loss and stability. As shown in Figure 19, QK-Norm is applied after the Query and Key projections, similar to output-LN. From another perspective, QK-Norm is also applied immediately before the attention calculation, akin to a Pre-LN approach. In our view, QK-Norm and Pre-LN (placed at $A^{2}$ and $A$ respectively in Figure 19) serve the same role but differ in certain details. As shown in Figures 20, 21, and 22, the two architectures exhibit comparable performance overall in terms of both training loss and stability.. However, Peri-LN provides a slight performance advantage over the OLMo2222-style Peri-LN in the $400$M- and $1$B-parameter models.
Figure 19: QK-layer normalization in the Attention module.
(a) Pre-training loss curve
(b) Gradient-norm curve
Figure 20: Comparison of pre-training loss and gradient norm between OLMo2-Style Peri-LN and the Peri-LN architecture. To ensure an accurate comparison, we present the pre-training loss over the final $5$B tokens. $400$M size model.
(a) Pre-training loss curve
(b) Gradient-norm curve
Figure 21: Comparison of pre-training loss and gradient norm between OLMo2-Style Peri-LN and the Peri-LN architecture. To ensure an accurate comparison, we present the pre-training loss over the final $5$B tokens. $1.5$B size model.
(a) Pre-training loss curve
(b) Gradient-norm curve
Figure 22: Comparison of pre-training loss and gradient norm between OLMo2-Style Peri-LN and the Peri-LN architecture. To ensure an accurate comparison, we present the pre-training loss over the final $5$B tokens. $3.2$B size model.
Appendix H Additional Details on Evaluation
H.1 Detailed Configurations
We utilized the Language Model Evaluation Harness [^6] https://github.com/EleutherAI/lm-evaluation-harnesswith the HuggingFace Transformers library (Gao et al., 2024; Wolf et al., 2020) to assess overall performance. We employ five different evaluation benchmarks: ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), Winogrande (Sakaguchi et al., 2021). During the pretraining stage, each model was trained under a controlled random seed. We used the training loss at iteration $14,000$—corresponding to the completion of $30$B tokens—as our main reference point. When calculating the evaluation score, diverged checkpoints were excluded.
H.2 Detailed Results on Benchmark Evaluations
In this section, we present the evaluation results for each model trained with five different training seeds. We exclude any diverged scores and average the remaining values, which are then reported in Table 2 in the main text.
H.2.1 Pre-Training
Table 6: Detailed results on pre-training the Peri-LN architecture. These results are averaged to produce the values reported in Table 2. SEED denotes pre-training seed.
Peri-LN SEED ARC-Easy HellaSwag PIQA SIQA Winogrande
1 0.5758 0.3803 0.6980 0.4115 0.5225
2 0.5728 0.3739 0.6915 0.4017 0.5367
400M 3 0.5842 0.3745 0.6986 0.4125 0.5249
4 0.5800 0.3722 0.6959 0.4038 0.5209
5 0.5627 0.3719 0.6899 0.4028 0.5320
1 0.6599 0.4437 0.7339 0.4304 0.5714
2 0.6591 0.4394 0.7399 0.4145 0.5699
1.5B 3 0.6625 0.4357 0.7372 0.4166 0.5627
4 0.6633 0.4367 0.7345 0.4222 0.5667
5 0.6637 0.4416 0.7361 0.4335 0.5612
1 0.6953 0.4734 0.7443 0.4417 0.5872
2 0.6839 0.4684 0.7427 0.4324 0.6054
3.2B 3 0.6902 0.4680 0.7486 0.4243 0.5967
4 0.6864 0.4700 0.7427 0.4273 0.5935
5 0.6806 0.4698 0.7372 0.4243 0.6054
Table 7: Detailed results on pre-training the Pre-LN architecture. These results are averaged to produce the values reported in Table 2. SEED denotes pre-training seed.
Pre-LN SEED ARC-Easy HellaSwag PIQA SIQA Winogrande
1 0.5669 0.3609 0.7008 0.4002 0.5359
2 Diverged
400M 3 0.5354 0.3328 0.6741 0.3905 0.4957
4 Diverged
5 0.5438 0.3314 0.6888 0.4012 0.4949
1 0.6326 0.4259 0.7242 0.4263 0.5691
2 0.6019 0.3924 0.7111 0.3992 0.5627
1.5B 3 0.6077 0.3932 0.7008 0.4125 0.5272
4 0.6111 0.3886 0.7187 0.4135 0.5225
5 0.6221 0.3941 0.7160 0.4099 0.5438
1 0.6688 0.4588 0.7470 0.4273 0.5919
2 Diverged
3.2B 3 Diverged
4 0.6359 0.4259 0.7301 0.4263 0.5564
5 Diverged
Table 8: Detailed results on pre-training the Post-LN architecture. These results are averaged to produce the values reported in Table 2. SEED denotes pre-training seed.
Post-LN SEED ARC-Easy HellaSwag PIQA SIQA Winogrande
1 0.3413 0.2881 0.6311 0.3378 0.5067
2 0.3691 0.2886 0.6132 0.3337 0.5099
400M 3 0.3632 0.2889 0.6257 0.3603 0.5051
4 0.3603 0.2920 0.6262 0.3490 0.5012
5 0.3510 0.2880 0.6170 0.3434 0.5209
1 0.4268 0.3121 0.6659 0.3628 0.5185
2 0.4196 0.3150 0.6654 0.3639 0.5004
1.5B 3 Diverged
4 0.4285 0.3212 0.6730 0.3511 0.4775
5 0.4419 0.3193 0.6643 0.3557 0.5154
1 0.4731 0.3427 0.6774 0.3664 0.5343
2 0.4638 0.3326 0.6779 0.3577 0.4917
3.2B 3 0.3956 0.3321 0.6143 0.3408 0.5067
4 0.4663 0.3380 0.6692 0.3685 0.5178
5 0.4663 0.3340 0.6839 0.3577 0.5043
H.2.2 Supervised Fine-Tuning
Table 9: Detailed results on SFT with Peri-LN architecture. These results are averaged to produce the values reported in Table 2. SEED denotes pre-training seed.
Peri-LN SEED ARC-Easy HellaSwag PIQA SIQA Winogrande
1 0.5800 0.3819 0.6991 0.4145 0.5328
2 0.5783 0.3783 0.6921 0.4038 0.5391
400M 3 0.5888 0.3806 0.6980 0.4222 0.5288
4 0.5892 0.3738 0.6948 0.4089 0.5099
5 0.5783 0.3757 0.6991 0.4099 0.5312
1 0.6633 0.4502 0.7356 0.4304 0.5746
2 0.6641 0.4437 0.7405 0.4207 0.5706
1.5B 3 0.6671 0.4454 0.7454 0.4207 0.5620
4 0.6700 0.4455 0.7378 0.4284 0.5659
5 0.6688 0.4478 0.7421 0.4324 0.5620
1 0.7058 0.4810 0.7486 0.4422 0.5880
2 0.6898 0.4774 0.7437 0.4391 0.6054
3.2B 3 0.6995 0.4770 0.7481 0.4278 0.5912
4 0.6911 0.4777 0.7432 0.4350 0.5943
5 0.6894 0.4781 0.7448 0.4319 0.6046
Table 10: Detailed results on SFT with Pre-LN architecture. These results are averaged to produce the values reported in Table 2. SEED denotes pre-training seed.
Pre-LN SEED ARC-Easy HellaSwag PIQA SIQA Winogrande
1 0.5762 0.3625 0.7078 0.4058 0.5343
2 N/A
400M 3 0.5370 0.3339 0.6757 0.3905 0.4972
4 N/A
5 0.5509 0.3372 0.6893 0.4074 0.4886
1 0.6385 0.4310 0.7247 0.4227 0.5620
2 0.6035 0.3934 0.7095 0.4038 0.5572
1.5B 3 0.6098 0.3944 0.7035 0.4150 0.5257
4 0.6208 0.3929 0.7182 0.4161 0.5272
5 0.6258 0.4017 0.7171 0.4181 0.5391
1 0.6785 0.4681 0.7568 0.4345 0.5825
2 N/A
3.2B 3 N/A
4 0.6427 0.4293 0.7274 0.4299 0.5580
5 N/A
Table 11: Detailed results on SFT with Post-LN architecture. These results are averaged to produce the values reported in Table 2. SEED denotes pre-training seed.
Post-LN SEED ARC-Easy HellaSwag PIQA SIQA Winogrande
1 0.4428 0.3307 0.6583 0.3797 0.5099
2 0.4280 0.3208 0.6404 0.3746 0.5178
400M 3 0.4693 0.3241 0.6578 0.3905 0.5122
4 0.4680 0.3247 0.6610 0.3726 0.4830
5 0.4520 0.3283 0.6572 0.3849 0.5225
1 0.5316 0.3774 0.6980 0.3889 0.5359
2 0.4731 0.3316 0.6719 0.3813 0.5028
1.5B 3 N/A
4 0.5387 0.3546 0.6779 0.3864 0.4909
5 0.5261 0.3510 0.6752 0.3767 0.5209
1 0.5623 0.4029 0.7008 0.3920 0.5051
2 0.5417 0.3644 0.6823 0.3833 0.5264
3.2B 3 0.4444 0.3604 0.6333 0.3618 0.5043
4 0.5400 0.3645 0.6844 0.3823 0.5020
5 0.5341 0.3677 0.6942 0.3976 0.5012

### v3

@@
-Peri-LN: Revisiting Layer Normalization in the Transformer Architecture
+Peri-LN: Revisiting Normalization Layer in the Transformer Architecture
 Abstract
-Designing Transformer architectures with the optimal layer normalization (LN) strategy that ensures large-scale training stability and expedite convergence has remained elusive, even in this era of large language models (LLMs).
-To this end, we present a comprehensive analytical foundation for understanding how different LN strategies influence training dynamics in large-scale Transformer training.
-Until recently, Pre-LN and Post-LN have long dominated standard practices despite their limitations in large-scale training.
-However, several open-source large-scale models have recently begun silently adopting a third strategy without much explanation.
-This strategy places layer normalization (LN) peripherally around sublayers, a design we term Peri-LN.
-While Peri-LN has demonstrated promising empirical performance, its precise mechanisms and benefits remain almost unexplored.
-Our in-depth analysis shows that Peri-LN strikes an ideal balance in variance growth—unlike Pre-LN and Post-LN, which are prone to vanishing gradients and “massive activations.” To validate our theoretical insight, we conduct large-scale experiments on Transformers up to $3.2$B parameters, showing that Peri-LN consistently achieves more balanced variance growth, steadier gradient flow, and convergence stability.
-Our results suggest that Peri-LN warrants broader consideration for large-scale Transformer architectures, providing renewed insights into the optimal placement and application of LN.
+Selecting a layer normalization (LN) strategy that stabilizes training and speeds convergence in Transformers remains difficult, even for today’s large language models (LLM).
+We present a comprehensive analytical foundation for understanding how different LN strategies influence training dynamics in large-scale Transformers.
+Until recently, Pre-LN and Post-LN have long dominated practices despite their limitations in large-scale training.
+However, several open-source models have recently begun silently adopting a third strategy without much explanation.
+This strategy places normalization layer peripherally around sublayers, a design we term Peri-LN.
+While Peri-LN has demonstrated promising performance, its precise mechanisms and benefits remain almost unexplored.
+Our in-depth analysis delineates the distinct behaviors of LN strategies, showing how each placement shapes activation variance and gradient propagation.
+To validate our theoretical insight, we conduct extensive experiments on Transformers up to $3.2$B parameters, showing that Peri-LN consistently achieves more balanced variance growth, steadier gradient flow, and convergence stability.
+Our results suggest that Peri-LN warrants broader consideration for large-scale Transformer architectures, providing renewed insights into the optimal placement of LN.
 normalization, transformers, pre-training, large language models, training stability
 1 Introduction
-Building on a rapidly expanding lineage of Transformer-based large language models, open-source models have shown remarkable impact (Hoffmann et al., 2022; Dubey et al., 2024; Guo et al., 2025).
+Building on a rapidly expanding lineage of Transformer-based large language models, open-source models have shown remarkable impact (Hoffmann et al., 2022; Guo et al., 2025; Yoo et al., 2024).
 As the demand for larger and more powerful models grows, various training stabilization techniques have been introduced (Yang et al., 2022; Zhai et al., 2023; Loshchilov et al., 2024).
-Among these, the choice of where and how to apply layer normalization (LN: LayerNorm or RMSNorm; Ba et al., 2016; Zhang & Sennrich, 2019) critically influences model convergence (Xiong et al., 2020; Kedia et al., 2024; Wortsman et al., 2024; Chung et al., 2024).
+Among these, the choice of where and how to apply layer normalization (LN: LayerNorm or RMSNorm; Ba et al., 2016; Zhang & Sennrich, 2019) critically influences model convergence (Xiong et al., 2020; Kedia et al., 2024; Wortsman et al., 2024).
 However, their immense computational requirements have restricted deeper exploration of the underlying Transformer structure.
@@
 Pre-LN (Dubey et al., 2024), by contrast, normalizes before passing the hidden state to the sub-layer (that is, $x+\mathrm{Module}(\mathrm{Norm}(x))$).
-While this can enhance gradient propagation, it also admits so-called “massive activations,” where hidden states grow exponentially across layers (Sun et al., 2024; Wortsman et al., 2024; Zhai et al., 2023).
+While this can enhance gradient propagation, it also admits so-called “massive activations,” where hidden states grow exponentially across layers (Sun et al., 2024).
 Previous studies on deep convolutional neural networks (CNNs) have analyzed the impact of batch normalization on variance changes during the initialization stage of ResNet architectures, demonstrating its relationship to model performance (De & Smith, 2020).
@@
 However, in the context of Transformer architectures, we observed that this variance growth at initialization does not persist as training progresses as shown in Figure 1.
-Section 3 and 4 provide a more detailed discussion of these hidden-state growth patterns.
+Sections 3, 4, 5, and 6 provide a more detailed discussion of these hidden-state growth patterns.
 Beyond these two common strategies, Post-LN and Pre-LN, a third LN placement has quietly emerged in large-scale open-source models: applying LN around the sub-layer, i.e., on both its input and output.
-Although recent open-source models (Rivière et al., 2024; OLMo et al., 2024) have quietly adopted such designs and demonstrated promising performance on a large scale, these efforts often appeared isolated, lacking a conceptual unifying framework or a thorough investigation into their benefits.
-In this paper, we coin the term Peri-LN111“Peri-” means “around,” reflecting that LN encapsulates the entire sub-layer.
-to unify these scattered approaches and highlight an underexplored avenue for stabilizing large-scale Transformer training.
-While prior work offers encouraging empirical results, there is still a lack of in-depth analysis explaining why and how Peri-LN can help mitigate issues that frequently arise with Post-LN and Pre-LN.
-By probing the forward and backward dynamics of Peri-LN, we provide fresh insights into where and when it may offer advantages over more widely adopted LN placements.
+Although recent open-source models (Team et al., 2025; Rivière et al., 2024; OLMo et al., 2024) have quietly adopted such designs and demonstrated promising performance on a large scale, these efforts often appeared isolated, lacking a conceptual unifying framework or a thorough investigation into their benefits.
+In this paper, we coin the term Peri-LN [^1] Peri- (Prefix) means “around,” reflecting that LN encapsulates the entire sub-layer.
+(e.g.
+peripherally) to unify these scattered approaches and highlight an underexplored avenue for stabilizing large-scale Transformer training.
+By dissecting the forward- and backward-pass dynamics of each LN strategy, we clarify how, when, and why they differ, interpreting these distinctions through the lens of training stability.
 Accordingly, this paper revisits LN placement in Transformers from both analytical and empirical perspectives. In particular, we:
   1. 1.
-    Present an in-depth empirical analysis of Post-LN and Pre-LN in large-scale Transformers, examining how variance and gradient properties evolve beyond initialization.
-By studying forward and backward propagation statistics, we clarify why Post-LN may suffer vanishing gradients and why Pre-LN can exhibit exponential growth of hidden state.
+    Present an in-depth analysis of Post-LN and Pre-LN in large-scale Transformers, examining how variance and gradient properties evolve beyond initialization.
   2. 2.
     Investigate Peri-LN to understand how normalizing both the inputs and outputs of each module moderates hidden-state behavior during forward and backward propagation, providing a systematic perspective on this underexplored alternative.
-We show that Peri-LN not only improves gradient stability and final loss but also plays a critical role in reducing hidden-state redundancy, as demonstrated by our ablation studies.
   3. 3.
-    Provide quantitative evidence on how large activation influences training stability, benchmark performance, and redundancy in hidden states.
-Drawing on theoretical bounds for gradient propagation as well as extensive experiments, we demonstrate that Peri-LN can mitigate excessive variance growth and gradient instability more effectively than commonly used alternatives.
+    Provide quantitative evidence on how large activation influences training stability, benchmark performance, and model behaviors.
 Figure 1: Illustration of hidden-state variance across different model depths and training iterations.
-From the initialization stage up to the point where 6.3 billion tokens were trained, we observed the variance growth of hidden states for Pre-LN and Post-LN architectures.
-The analysis was conducted using a $1.5$B-parameter model, and consistent trends were observed across models of different sizes.
-Detailed settings and more results are in Section 4.4.2.
+From initialization through training on $6.3$ billion tokens, we observe the growth in hidden-state variance for both Pre-LN and Post-LN architectures.
+The analysis is based on a $1.5$B-parameter model.
+Detailed settings and additional results are provided in Section 5.1.
 2 Background and Motivation
 The analysis of activation variance at model initialization has long been central to understanding normalization layers and enhancing stability in convolutional neural networks (CNNs) (De & Smith, 2020; He et al., 2016; Brock et al., 2021a).
-De & Smith (2020) showed that batch normalization in residual blocks can bias networks toward the identity function, thereby stabilizing gradients and improving overall training dynamics.
-Similar investigations have emerged for Transformer architectures, examining how variance propagates and how gradients behave in both post-layer normalization (Post-LN) (Vaswani et al., 2017) and pre-layer normalization (Pre-LN) (Dubey et al., 2024) configurations (Xiong et al., 2020; Kedia et al., 2024; Wortsman et al., 2024; Li et al., 2024b).
-Early work comparing Post- and Pre-LN primarily focused on gradient scales and loss behavior.
+Specifically, De & Smith (2020) showed that batch normalization in residual blocks can bias networks toward the identity function, thereby stabilizing gradients and improving overall training dynamics.
+Similar investigations have emerged for Transformer architectures, examining how variance propagates and how gradients behave in both post-layer normalization (Post-LN) and pre-layer normalization (Pre-LN) configurations (Xiong et al., 2020; Kedia et al., 2024; Wortsman et al., 2024).
+Early work comparing Post- and Pre-LN primarily focused only on the initialization stage.
 Xiong et al.
 (2020) observed that Pre-LN architectures tend to exhibit more stable gradients, but can still encounter issues such as gradient spikes and divergence, especially in deeper models or large-scale pre-training scenarios (Zhai et al., 2023; Wortsman et al., 2024; Fishman et al., 2024; Chung et al., 2024).
-Among these challenges, the phenomenon of “massive activations” has attracted particular attention (Dettmers et al., 2022; Yu et al., 2024; Fishman et al., 2024).
-Sun et al.
+Among these challenges, the phenomenon of “massive activations” has attracted attention (Dettmers et al., 2022; Yu et al., 2024; Fishman et al., 2024).
+Notably, Sun et al.
 (2024) identified that in Pre-LN architectures, large spikes in activation magnitude can persist across layers due to residual connections.
 These massive activations act as fixed biases, potentially narrowing the model’s focus to certain tokens and may influence generalization.
-However, the underlying mechanisms behind these large values—and their exact impact on the training process—remain not yet well understood.
+However, the underlying mechanisms behind these large values, and their exact impact on the training process, remain not yet well understood.
 Analytical work has provided theoretical frameworks to explain phenomena like gradient explosion and vanishing in Transformers.
@@
 Recent studies have discussed how Pre-LN architectures can allow large values from Attention or MLP modules to flow unimpeded through residual connections (Csordás et al., 2024; Fishman et al., 2024; Zhai et al., 2023; Wortsman et al., 2024), but the precise impact of this behavior on large-scale training remains insufficiently explored.
-These observations underscore the ongoing need to clarify how activation dynamics, normalization strategies, and architectural choices interact, especially in large-scale models.
-In response, this work aims to deepen our understanding of activation evolution during Transformer training under different normalization architectures, focusing on the role of massive activations and their effects on overall stability and performance.
-We defer an extended discussion of the related literature to Appendix A, owing to space limitations.
-3 Comparative Analysis
-In this section, we discuss how different placements of layer normalization (LN  [^2] Unless stated otherwise, LN refers to both LayerNorm and RMSNorm.) in Transformer architecture affect both training stability and the statistics of hidden states (activations  [^3] We use “hidden state” and “activation” interchangeably.).
+These observations underscore the ongoing need to clarify how activation dynamics and normalization strategies interact, especially in large-scale training.
+In response, this work aims to deepen our understanding of how normalization strategies influence Transformer training, with particular attention to the emergence of large activations and their implications for stability and performance.
+3 Normalization Strategies
+In this section, we discuss how different placements of layer normalization (LN  [^2] Unless stated otherwise, LN refers to both LayerNorm (Ba et al., 2016) and RMSNorm (Zhang & Sennrich, 2019).) in Transformer architecture affect both training stability and the statistics of hidden states (activations  [^3] We use “hidden state” and “activation” interchangeably.).
 3.1 Post- & Pre-Normalization in Transformers
@@
 Figure 2: Placement of normalization in Transformer sub-layer.
-Table 1: Intuitive comparison of normalization strategies.
-Strategy Variance Growth Gradient Stability
-Post-LN Mostly constant Potential for vanishing
-Pre-LN Exponential in depth Potential for explosion
-[…]
 3.2 Variance Behavior from Initialization to Training
@@
 To bridge the gap from initialization to the fully trained stage, we extend our empirical observations in Figure 1 beyond initial conditions by tracking how these variance trends evolve at intermediate points in training.
-We find that Post-LN maintains a roughly constant variance, which helps avert exploding activations.
-Yet as models grow deeper and training proceeds, consistently normalizing $x_{l}+\mathrm{Module}(x_{l})$ can weaken gradient flow, occasionally causing partial vanishing gradients and slower convergence.
+During training, we find that Post-LN maintains a roughly constant variance, which helps avert exploding activations.
+However, as models grow deeper and training progresses, consistently normalizing $x_{l}+\mathrm{Module}(x_{l})$ can weaken gradient flow, occasionally causing partial vanishing gradients and slower convergence.
 In contrast, Pre-LN normalizes $x_{l}$ before the module but leaves the module output unnormalized, allowing hidden-state variance to accumulate exponentially once parameter updates amplify the input.
 Although Pre-LN preserves gradients more effectively in earlier stages, this exponential growth in variance can lead to “massive activations” (Sun et al., 2024), risking numeric overflow and destabilizing large-scale training.
-We reconfirm this in Section 4.
-Takeaways.
-  • Keeping the Highway Clean: Post-LN’s Potential for Gradient Vanishing and Slow Convergence.
-When layer normalization is placed directly on the main path (Placement $C$ in Figure 2), it can cause gradient vanishing and introduce fluctuations in the gradient scale, potentially leading to instability.
-  • Maintaining a Stable Highway: Pre-LN May Not Suffice for Training Stability.
+Takeaways from Pre-LN & Post-LN.
+(1) Keeping the Highway Clean: Post-LN’s Potential for Gradient Vanishing and Slow Convergence.
+When layer normalization is placed directly on the main path (Placement $C$ in Figure 2), it can cause gradient vanishing and introduce fluctuations in the gradient scale, potentially leading to instability (Xiong et al., 2020).
+(2) Maintaining a Stable Highway: Pre-LN May Not Suffice for Training Stability.
 Pre-LN does not normalize the main path of the hidden states, thereby avoiding the issues that Post-LN encounters.
 Nevertheless, a structural characteristic of Pre-LN is that any large values arising in the Attention or MLP modules persist through the residual identity path.
-In particular, as shown in Figure 1, the exponentially growing magnitude and variance of the hidden states in the forward path may lead to numerical instability and imbalance during training.
-Recent open-sourced Transformer architectures have adopted normalization layers in unconventional placements.
-Models like Gemma2222 and OLMo2222 utilize normalization layers at the module output (Output-LN), but the benefits of these techniques remain unclear (Rivière et al., 2024; OLMo et al., 2024).
-To investigate the impact of adding an Output-LN, we explore the peri-layer normalization architecture.
-3.3 Placing Module Output Normalization
+In particular, as shown in Figure 1, the exponentially growing magnitude and variance of the hidden states in the forward path may lead to numerical instability.
+3.3 Peri-Normalization in Transformers
+Recent open-source Transformer architectures have placed normalization layers in unconventional placements (Rivière et al., 2024; Team et al., 2025; OLMo et al., 2024).
+In particular, these models apply an additional normalization layer at the module output (Output-LN), yet the benefits of this design choice remain unclear.
+To assess the impact of Output-LN, we analyze the Peri-LN architecture.
 Peri-LN.
@@
   1. 1.
-    Initial Embedding Normalization:
+    (Optional) Initial Embedding Normalization:
     $y_{o}=\mathrm{Norm}(x_{o}),$
@@
     Input- & Output-Normalization per Layer:
-    $y_{l}=x_{l}+\mathrm{Norm}\Bigl{(}\mathrm{Module}\bigl{(}\mathrm{Norm}(x_{l})\bigr{)}\Bigr{)},$
+    $y_{l}=x_{l}+\mathrm{Norm}\Bigl{(}\mathrm{Module}\bigl{(}\mathrm{Norm}(x_{l})\bigr{)}\Bigr{)},$ (3)
   3. 3.
@@
 For clarity, the locations of normalization layers in the Post-, Pre-, and Peri-LN architectures are illustrated in Figure 2.
+Both the latest Gemma (Team et al., 2025; Rivière et al., 2024) and OLMo (OLMo et al., 2024) model families, which apply output layer normalization, adopt the same peri-normalization strategy.
+However, neither line of work rigorously examines how this placement constrains variance or mitigates large residual activations.
+Our study extends these open-sourced large-scale models by providing both theoretical and empirical insights into the Peri-LN scheme.
 Controlling Variance & Preserving Gradients.
-By normalizing both the input and output of each sub-layer, Peri-LN constrains the residual spikes common in Pre-LN, while retaining a stronger gradient pathway than Post-LN.
-Concretely, if $\mathrm{Norm}(\mathrm{Module}(\mathrm{Norm}(x_{l})))$ has near-constant variance $\beta_{0}$, then
-$\mathrm{Var}(x_{l+1})\;\approx\;\mathrm{Var}(x_{l})+\beta_{0},$
-leading to linear or sub-exponential hidden state growth rather than exponential blow-up. We empirically verify this effect in Section 4.4.2.
+By normalizing both the input and output of each sub-layer, Peri-LN constrains the residual spikes commonly observed in Pre-LN, while maintaining a stronger gradient pathway than Post-LN.
+Concretely, if $\mathrm{Norm}(\mathrm{Module}(\mathrm{Norm}(x_{l})))$ exhibits near-constant variance $\beta_{0}$, then
+$\mathrm{Var}(x_{l+1})\;\approx\;\mathrm{Var}(x_{l})+\beta_{0},$ (4)
+resulting in linear or sub-exponential growth of activations, in contrast to the exponential growth patterns of Pre-LN.
+Although Pre-LN and Peri-LN exhibit comparable, roughly linear variance growth at initialization (De & Smith, 2020; Xie et al., 2023), their trajectories diverge once training begins.
+The additional normalization layer (Output-LN) in Peri-LN preserves the conditions of Eq.
+4, enabling the model’s hidden states to remain better conditioned.
+By contrast, the rapid surge in variance observed in Pre-LN can trigger instability during the early stages of training, an effect we quantify empirically in Sections 5 and 6.
 (a) Learning rate exploration
@@
 (c) Gradient-norm
-Figure 3: Performance comparison of Post-LN, Pre-LN, and Peri-LN Transformers during pre-training.
+Figure 3: Performance comparison of Post-, Pre-, and Peri-LN Transformers during pre-training.
 Figure 3(a) llustrates the pre-training loss across learning rates.
-Pre-training loss and gradient norm of best performing $400$M size Transformers are in Figure 3(b) and 3(c).
-Consistent trends were observed across models of different sizes.
-(a) Training loss
-(b) Loss in the final $5$B token interval
-(c) Gradient-norm
-Figure 4: Freezing learnable parameter $\gamma$ of output normalization layer in Peri-LN. we set $\gamma$ to its initial value of $1$ and keep it fixed.
-Open-Sourced Peri-LN Models: Gemma2222 & OLMo2222.
-Both Gemma2222 and OLMo2222, which apply output layer normalization, employ the same peri-normalization strategy within each Transformer layer.
-However, neither model rigorously examines how this placement constrains variance or mitigates large residual activations.
-Our work extends Gemma2222 and OLMo2222 by offering both theoretical and empirical perspectives within the Peri-LN scheme.
-Further discussion of the OLMo2222 is provided in Appendix G.
+Pre-training loss and gradient norm of best performing $400$M size Transformers are in Figure 3(b) & 3(c).
+(a) Divergence at seed $2$
+(b) Loss spike at seed $3$
+(c) Gradient spikes at seed $5$
+(d) Loss spikes at seed $5$
+Figure 4: Common case of early stage instability in pre-training.
+In most of our experiments across different random seeds, the Pre-LN architecture exhibited early-stage instability.
+Although we initially suspected that a high learning rate might be the root cause, lowering it did not substantially mitigate these issues.
+By contrast, under the same settings, Peri-LN displayed stable training curves.
 3.4 Stability Analysis in Normalization Strategies
-We analyze training stability in terms of the magnitude of activation.
-To this end, we examine the gradient norm with respect to the weight of the final layer in the presence of massive activation.
-For the formal statements and detailed proofs, refer to Appendix D.
+Xiong et al.
+(2020) showed that, at initialization, Pre-LN exhibits smaller gradient scales at the final layer compared to Post-LN, with respect to model depth.
+In this study, we broaden our analysis beyond initialization to monitor hidden state variance over the full course of training.
+Building on the earlier observation that the deepest layer exhibits the largest activations, we focus on this surge in the final layer under the Peri-LN strategy to clarify its impact on training stability.
+To this end, we analyze stability by examining the gradient norm with respect to the final layer weights in the presence of massive activation.
+Formal statements and detailed proofs are presented in Appendix C.
 Proposition 3.1 (Informal).
@@
 In this case, even when a massive activation $\|h\|$ occurs, $\mathrm{Norm}(\cdot)$ introduces a damping factor $\|a\|$, which ensures that the gradient norm $\|\partial\mathcal{L}/\partial W^{(2)}\|$ remains bounded.
-(3) Post-LN (vanishing gradient). Consider the following sequence of operations:
-$a=\mathrm{MLP}(x),o=x+a,\tilde{o}=\mathrm{Norm}(o),$ (7)
-then
-$\left\lVert\frac{\partial\mathcal{L}(\tilde{o})}{\partial W_{i,j}^{(2)}}\right\rVert\;\leq\;\frac{4\,\gamma\,\sqrt{D}\,\|h\|}{\|x+a\|},$ (8)
-where $h:=\mathrm{ReLU}\left(xW^{(1)}+b^{(1)}\right)$.
-In this case, when a massive activation $\|h\|$ occurs, $\mathrm{Norm}(\cdot)$ introduces an overly suppressing factor $\|x+a\|$, which contains a separate huge residual signal $x$, potentially leading to a vanishing gradient $\|\partial\mathcal{L}/\partial W^{(2)}\|$.
-We have compiled a Table 1 that provides a overview of the variance and gradient intuition for each layer normalization strategy.
+The layer-wise amplification documented in §3.2, combined with the bounds in Proposition 3.1, naturally explains the gradient spikes, and occasional divergences that arise in Pre-LN during large-scale pre-training.
+We revisit this phenomenon in §4.3.
+By contrast, the additional normalization in Peri-LN acts as a self-regularizing mechanism that damps variance growth, making the architecture less sensitive to large activations and therefore more stable in practice.
+The formal analysis for Post-LN is deferred to Appendix B.
+4 Experiments
+In this section, we provide a comprehensive comparison of Post-, Pre-, and Peri-Layer Normalization (LN) across large-scale Transformer pre-training, instruction-tuning, and subsequent evaluations on the language domain.
 Table 1: Average benchmark scores (with standard deviations) across $5$ different training seeds for Post-, Pre-, and Peri-Layer Normalization language models.
 Each model size excludes the embedding parameters.
-Loss denotes the evaluation loss on random samples of the C4444 dataset (Raffel et al., 2020).
+Loss denotes the evaluation loss on random samples of the C4444 dataset.
 Arch.
@@
 denotes the averaged benchmark score across tasks of instruction fine-tuned models.
-When calculating the evaluation score, diverged checkpoints were excluded.
+Diverged checkpoints are excluded from the evaluation score computation.
 Size Arch. ARC-Easy HellaSwag PIQA SIQA Winogrande Avg. $\uparrow$ Loss $\downarrow$ SFT Avg. $\uparrow$
@@
 Peri-LN 68.73 ±$\pm 0.57$ 46.99 ±$\pm 0.21$ 74.31 ±$\pm 0.41$ 43.00 ±$\pm 0.73$ 59.76 ±$\pm 0.78$ 58.56 3.11 59.02
-4 Experiments
-In this section, we provide a comprehensive empirical comparison of Post-, Pre-, and Peri-Layer Normalization (LN) across large-scale Transformer pre-training and subsequent evaluations on the language domain.
 4.1 Experimental Setting
-Excluding the embedding parameters, model sizes were set to $400$M, $1.5$B, and $3.2$B parameters, respectively.
-Each model was trained on $30$ billion tokens.
-For a fair comparison, we pre-trained each model with five different training seeds.
-We conducted a systematic exploration of learning rates, ranging from $1\times 10^{-4}$ to $5\times 10^{-3}$.
-The global batch size is set to $256$.
-We use Adam optimizer with cosine learning rate scheduler.
-The sequence length was set to $8192$, and the weight decay coefficient was fixed at $0.033$.
-We employed Megatron-LM [^4] https://github.com/NVIDIA/Megatron-LM to pre-train the Transformers under each layer normalization strategy.
-We used the DCLM-baseline dataset (Li et al., 2024a) for our experiments, along with the “cl100100100100k_base” version of the TikToken tokenizer [^5] https://github.com/openai/tiktoken.
-Unless otherwise noted, most training and model settings followed those of the DCLM experiments.
-For layer normalization, we primarily employed RMSNorm.
-Further details are provided in Appendix B.
+Excluding the embedding parameters, the model size is set to the parameters $400$M, $1.5$B and $3.2$B, respectively.
+Each model is trained on $30$ billion tokens.
+To ensure reliable validation, we pre-train each model with five different training seeds in all experiments.
+We perform a exploration of the learning rates, ranging from $1\times 10^{-4}$ to $5\times 10^{-3}$ to identify the U-shaped pattern for each LN strategy.
+The sequence length is set to $8192$, and the weight decay coefficient is fixed at $0.033$.
+We employ Megatron-LM [^4] https://github.com/NVIDIA/Megatron-LM to pre-train the Transformers under each LN strategy.
+We use the DCLM-baseline dataset (Li et al., 2024a), along with the “cl100100100100k_base” version of the TikToken tokenizer [^5] https://github.com/openai/tiktoken.
+Unless otherwise noted, most training and model configurations follow those of the DCLM experiments(Li et al., 2024a).
+For normalization layer, we primarily employ RMSNorm.
+Further details are in Appendix D.
 4.2 Pre-Training Large Language Models
@@
 Since Pre-LN shows best performance at learning rate $2\times 10^{-3}$ across all model size, we set this to the default learning rate for Pre-LN and Peri-LN.
-Unlike Pre-LN, Post-LN’s appropriate learning rate lies in a lower range, so we have provided a separate summary in Appendix C.1.
+Unlike Pre-LN, Post-LN’s appropriate learning rate lies in a lower range, so we provide a separate summary in Appendix E.1.
 In Figures 3(b) and 3(c), we compare the pre-training loss and the gradient norm curve at each LN strategy’s best-performing learning rate of $400$M size models.
-The same trend was observed across different model sizes.
-In particular, when sweeping over training seeds, Pre-LN and Post-LN exhibited many spikes in the gradient norm curve, whereas Peri-LN showed relatively few spikes, supporting Proposition 3.1.
-This reduction is consistent across both large and small learning rates.
-We provide additional results in Appendix F, including experiments using LayerNorm (instead of RMSNorm), analyses of varying sequence lengths, the effect of different training-token counts, and an ablation study on embedding LN.
-We also investigate pre-training using the OLMo2222-style Peri-LN architecture in Appendix G.
-These findings are likewise consistent with those previously reported.
-(a) Absolute magnitude growth
-(b) Variance growth
-Figure 5: This figure shows the forward growth patterns of hidden states for different architectures, highlighting the structural impact of normalization placement.
-Each model contains $1.5$ billion parameters (excluding the embedding size).
-We confirmed that the observed trend remains consistent across all model sizes.
-(a) Grad-norm at init
-(b) Grad-norm at final
-(c) Grad-variance at init
-(d) Grad-variance at final
-Figure 6: Comparison of the backward gradient norm and variance for Post-LN, Pre-LN, and Peri-LN Transformers at initialization and at the final stage.
-Model size is $1.5$B.
-We confirmed that the observed trend remained consistent across all model sizes.
-init denotes the initialization of the model parameters.
+The same trend is observed across different model sizes (§E).
+In particular, when we sweep over training seeds and learning rates, Pre-LN frequently exhibits spikes in the gradient-norm curve, whereas Peri-LN shows comparatively few, thereby supporting Proposition 3.1.
+(a) Gradient-norm at seed $5$
+(b) Gradient-norm at seed $4$
+Figure 5: Final-layer gradient norms for seeds $4$ and $5$.
+4.3 Early Stage Instability in Pre-Training
+Early in pre-training, Pre-LN models consistently show gradient spikes, loss surges, and occasional divergence across seeds and scales (Fig.
+4).
+These issues are far less pronounced in Peri-LN.
+We posit that the instability of Pre-LN arises from three factors: (1) the hidden state variance exhibits a sudden surge from initialization through the early stages of optimization, deviating from the linear trend predicted by Eq.
+4 (see §3.3); (2) the exponential growth of hidden state variance across both depth and training steps; and (3) the instability caused by the massive activations (Proposition 3.1).
+Among these, we highlight the variance growth along the main path as the principal driver of the observed divergence.
+To corroborate this, Section 6 presents targeted experiments that manipulate weight decay and weight initialization schemes, demonstrating how curbing extreme variance mitigates the instability of each LN strategy.
+The curves in 4(a), 4(b), and 4(c) are from a $400$M model, whereas 4(d) corresponds to a $1.5$B model.
+4.4 Gradient Norm of the Final-layer
+Motivated by Proposition 3.1, we track the final-layer gradient-norm in two representative runs selected from five training seeds.
+Figure 5(a), now including the newly added Peri-LN results, confirms the hierarchy reported by Xiong et al.
+(2020): whenever training remains stable, the gradient norms satisfy Post-LN $>$ Pre-LN $>$ Peri-LN.
+However, Figure 5(b) shows a run in which the Pre-LN model diverges even though every hyperparameter matches the stable run except for the random seed.
+In Section 5.1, we examine this failure in greater depth and relate it to Proposition 3.1.
+The curves in Figure 5 are obtained from $400$M models.
 4.5 Benchmark Evaluations & Instruction Tuning
-To evaluate how well the pre-trained model’s training loss aligns with its benchmark performance, we conducted five separate benchmark evaluations.
-Furthermore, to examine how supervised fine-tuning (SFT) improves scores and instruction-following capabilities under different layer normalization strategies in Transformer architectures, we performed additional training using the LIMA dataset (Ouyang et al., 2022; Zhou et al., 2023).
-Diverged checkpoints were excluded when calculating the evaluation score (mostly occurs in Pre-LN).
-Additional training hyperparameters for SFT are given in Appendix B.2.
+To evaluate how well the pre-training loss aligns with its benchmark performance, we conduct five separate benchmark evaluations.
+Furthermore, to investigate instruction-following capabilities under different layer normalization strategies, we conduct additional training using the LIMA dataset (Ouyang et al., 2022; Zhou et al., 2023).
+Diverged checkpoints are excluded from the evaluation score computation (mostly occurs in Pre-LN).
+Additional training hyperparameters for SFT are given in Appendix D.2.
 As shown in Table 1, Peri-LN consistently demonstrates superior performance across all model sizes.
@@
 For the evaluation loss, we used $10$K random samples from the C4444 dataset (Raffel et al., 2020).
-Detailed configurations and individual scores are provided in Appendix H.
-4.4 Systematic Analysis
-Despite emerging empirical evidence that Peri-LN can stabilize training and bolster performance, many open questions remain: Which design factors are crucial for reaping its benefits at scale (§4.4.1)?
-How does it regulate hidden-state variance and gradient flow more effectively (§4.4.2, §4.4.3)?
-And why might certain variants outperform standard LN placements in large Transformer architectures (§4.4.4)?
-Addressing these points, we present a systematic analysis of Peri-LN’s mechanics at following sub-sections.
-4.4.1 Learnable Parameter $\gamma$ of RMSNorm
-To investigate the impact of module output normalization on training stability, as proposed in the Proposition 3.1, we fixed the learnable parameter $\gamma$ of RMSNorm to $1$, isolating the effect of normalization.
-As shown in Figure 4, we observe that simply normalizing the output of each module reduces gradient spikes and provides improvements in the loss comparing to the pre-normalization only transformers.
-Nonetheless, we also confirmed that allowing $\gamma$ to be learnable yields slightly better performance.
-We confirm that this trend remains consistent in both the $400$M and $1.5$B models.
-In this experiment, we omitted Peri-LN’s embedding layer normalization in order to isolate and evaluate the precise role and benefits of output layer normalization.
-(a) At initialization
+Detailed settings and individual benchmark scores are provided in Appendix J.
+(a) Absolute magnitude growth
+(b) Variance growth
+Figure 6: Forward hidden state growth patterns for each LN strategy in a $1.5$B-parameter Transformer.
+(a) Grad-norm at init.
+(b) Grad-norm at final
+(c) Grad-variance at init.
+(d) Grad-variance at final
+Figure 7: Backward gradient norm and variance of $1.5$B Post-, Pre-, and Peri-LN Transformers at initialization (init.) and final training.
+5 Analysis
+Despite emerging evidence that Peri-LN outperforms Post- and Pre-LN, key uncertainties remain: How do different LN placements shape hidden-state statistics and gradient flow (§5.1, §5.2)?
+What role does the Output-LN scale parameter $\gamma$ play (§5.3)?
+And why does Peri-LN produce more distinctive representations than its counterparts (§5.4)?
+The subsections below tackle these questions in turn.
+5.1 Growth of Hidden State
+To examine in greater depth how Peri-LN affects forward propagation, we analyze the absolute magnitude and variance of the hidden states using $1,000$ samples from the Wikitext dataset (Merity et al., 2016).
+Figure 6 shows how different normalization strategies influence forward-path hidden states over the course of training and across model depth.
+We observe the same pattern across all models trained with five different random seeds (§F).
+Across layers, Post-LN maintains stable hidden state magnitudes and variances because the main path includes a normalization layer.
+In contrast, Pre-LN omits normalization after each attention and MLP sub-layer, so the magnitude and variance of the hidden states grow exponentially after the residual addition.
+For Peri-LN, which adds an Output-LN, these statistics remain comparatively well controlled.
+Across training iterations, Post-LN’s block-level normalization continues to suppress large shifts, preventing substantial drift in magnitude or variance.
+Pre-LN starts with an approximately linear variance profile at initialization but escalates exponentially to extremely large values as optimization proceeds.
+Peri-LN again exhibits only moderate fluctuations, owing to Output-LN’s consistent regulation of hidden-state statistics.
+Further discussion appears in Section 6.
+(a) Training loss
+(b) Loss in the final $5$B token interval
+(c) Gradient-norm
+Figure 8: Freezing learnable parameter $\gamma$ of output normalization layer in Peri-LN. we set $\gamma$ to its initial value of $1$ and keep it fixed.
+(a) After $30$B tokens training
 (b) Learnable scale $\gamma$ in Output-LN
-(c) After $30$B tokens training
-(d) Frozen $\gamma$ in Output-LN
-Figure 7: Angular distance of hidden state is presented in Figure  7(a)7(c), and 7(d).
-In Figure 7(b), we monitor $\gamma$ of every Output-LN in Peri-LN during training.
-We use $30$B tokens trained $400$M size model in this experiments.
-4.4.2 Growth of Hidden State
-To examine in greater depth how Peri-LN affects the forward propagation, we analyzed the absolute magnitude and variance of the hidden states using $1,000$ samples from the Wikitext dataset (Merity et al., 2016).
-Figure 5 illustrates how normalization strategies influence the forward-path hidden states over the course of training.
-Two particular aspects of hidden-state behavior were noted: First, with respect to model depth, Post-LN maintains stable hidden-state magnitude and variance owing to the presence of a normalization layer in the main path (“Highway”).
-By contrast, because Pre-LN does not normalize the outputs of each Attention and MLP module, the magnitude and variance of the hidden states exhibit exponential growth following the addition operation in the residual path.
-For the Peri-LN architecture, which leverages Output-LN, we observed that hidden states remain comparatively well-managed.
-Next, regarding changes over training iterations, Post-LN continues to normalize each block’s output, which prevents any substantial increase or decrease in the trend as training progresses.
-Meanwhile, Pre-LN exhibits a relatively linear variance distribution at initialization, but escalates exponentially to extremely large values over successive iterations.
-In contrast, Peri-LN fluctuates more moderately, owing to the role of Output-LN in controlling hidden-state magnitude and variance throughout training.
-4.4.3 Layer-wise Gradient Norm & Variance
+Figure 9: Angular distance between hidden states after training. Fig. 9(b) monitor $\gamma$ of every Output-LN in Peri-LN during training.
+5.2 Layer-wise Gradient Norm & Variance
 Ensuring a uniform gradient flow in large-scale model training is crucial for balanced learning across the entire network (Yang & Hu, 2021; Yang et al., 2024).
@@
 This stability is particularly beneficial in deeper architectures, where balanced gradient flow is essential for effective backpropagation.
-4.4.4 Hidden State Representation
-We utilize angular distance (Li et al., 2024b) as a metric to assess redundancy between hidden states at initialization and after training.
-This approach allows us to quantify how similar or distinct the representations were across layers.
-As shown in Figure 7(a), Post-LN exhibits smaller angular distances due to the LN being located on the main path, whereas Pre-LN and Peri-LN begin with very similar states.
-As shown in Figure 7(c), at the end of training, Pre-LN tends to produce more redundant hidden state representations compared to the others.
-This phenomenon may stem from Pre-LN’s repeated residual additions, which amplify certain representations over others.
-To investigate this further, we focus on module output normalization, which is the main distinguishing factor between Pre-LN and Peri-LN.
-As shown in Figure 7(b), the learnable scale starts around $1$ in the early stages of training and gradually changes with increasing depth.
-Because Peri-LN preserves the identity path, it appears to adjust the scale of the module output accordingly.
-This suggests that the exponential growth of the main path’s magnitude in Pre-LN diminishes the relative contribution of individual modules, resulting in more redundant hidden representations.
-To support this, Figure 7(d) shows that fixing the learnable scale of Peri-LN’s module output LN at $1$ causes the main path contribution to decrease in deeper layers.
-This finding confirms the role of module output normalization in controlling hidden state redundancy.
+5.3 Learnable Parameter $\gamma$ of RMSNorm
+To investigate the impact of module output normalization on training stability, as proposed in the Proposition 3.1, we fix the learnable parameter $\gamma$ of RMSNorm to $1$, isolating the effect of normalization.
+As illustrated in Figure 8, adding output normalization to each sub-layer suppresses gradient spikes and lowers the loss relative to Transformers that employ only pre-normalization.
+Nonetheless, we also confirm that allowing $\gamma$ to be learnable yields slightly better performance.
+The trend persists consistently across model scales and random seeds.
+In this experiment, we omit Peri-LN’s embedding layer normalization in order to isolate and evaluate the precise role and benefits of output-LN.
+5.4 Hidden State Representation
+To assess hidden state redundancy after training, we employ angular distance (Li et al., 2024b), which quantifies how similar or distinct the layer representations are.
+As Figure 9(a) illustrates, Pre-LN produces markedly more redundant hidden states than the other variants by the end of training.
+We attribute this effect to the exponential growth of the main residual path in Pre-LN, which diminishes the relative contribution of individual sub-layers.
+In contrast, Peri-LN retains an identity path whose learnable scale begins near 1 and gradually adjusts with depth (Figure 9(b)), thereby moderating redundancy.
+These observations highlight the role of module-output normalization in controlling hidden state similarity.
+All statistics are computed on $256$ random samples from RedPajama-Data-1T (Computer, 2023).
+Appendix L includes additional figures and initialization comparisons.
+6 Ablation Study
+(a) Variance growth by weight decay
+(b) Training loss by weight decay
+(c) Variance growth by weight init.
+Figure 10: Effects of weight decay and initialization (init.) on massive activations.
+10(a) Strong weight decay relieves the variance explosion in Pre-LN.
+10(b) Strong weight decay ($0.33$, which is 10×10$\times$10 × the baseline.) suppresses Pre-LN divergence.
+10(c) Smaller-scale initialization slightly curbs Pre-LN variance, while Peri-LN remains bounded regardless.
+$d$ denotes the model’s hidden dimensionality.
+To probe massive activations across conditions, we sweep weight decay coefficient and initialization variance for both Pre- and Peri-LN models, holding other settings fixed.
+Per-run results and detailed settings are in Appendix G.8.
+6.1 Weight Decay
+In Figure 10(a), stronger L2222 regularization markedly lowers the variance curve, confirming that heavier weight decay directly curbs forward-path explosions in Pre-LN.
+In contrast, the same increase in weight decay reduces Peri-LN’s variance growth only marginally.
+We take the stable run initialized with seed $3$ (Table 11) and sweep the weight decay coefficient.
+Table 7 in the Appendix further shows that, irrespective of the presence of massive activations, Peri-LN achieves better performance than Pre-LN.
+To further probe stability under varying degrees of massive activation, we replicate the previously divergent run (seed $4$) and repeat the same weight decay sweep.
+As Figure 10(b) demonstrates, raising the weight decay coefficient from the baseline $0.033$ to $0.33$ (a tenfold rise) prevents divergence, providing empirical support for Proposition 3.1.
+Nevertheless, strong weight decay can stabilize Pre-LN, it still fails to close the performance gap relative to Peri-LN.
+6.2 Weight Initialization
+As the initialization variance increases, the severity of massive activations rises correspondingly for Pre-LN (Figure 10(c)); at the largest variance, the model diverges outright (Appendix G.8.3).
+Pre-LN therefore displays marked sensitivity to its initial conditions.
+In contrast, across the same range of ablations, Peri-LN’s loss curves and activation variances shift only marginally.
+This hyperparameter-insensitive robustness is corroborated by the low downstream standard deviations reported in Table 1.
+6.3 Additional Results
+For brevity, we defer an extensive set of supplementary experiments to the appendix.
+Appendix G reports the core robustness checks: replacing RMSNorm with LayerNorm, varying sequence lengths, reducing pre-training budgets, and ablating embedding-level normalization.
+OLMo2222-style Peri-LN pre-training runs appear in Appendix H.
+Stochastic gradient descent (SGD) baselines are summarized in Appendix I.
+Alternative LN placements, extending Figure 2, are provided in Appendix G.6.
+Across all settings, the results are consistent with the trends presented in the main Section 4 and 5, further substantiating our conclusions.
+7 Implications
+This section integrates our findings into practical guidance on variance-driven stability and precision constraints in large-scale Transformers.
+7.1 Mitigating Variance-Driven Instability via Peri-LN
+Pre-LN, the prevailing normalization strategy, is inherently prone to unchecked growth in activation variance (§5), which in turn induces numerical instability during training.
+Our extensive empirical analysis shows that Peri-LN— which normalizes the outputs of the Attention and MLP sub-layers—markedly curbs this variance and often prevents divergence (§4 & §6).
+Proposition 3.1 formalizes how excessive variance amplifies gradient norms, clarifying its causal role in destabilizing large-scale pre-training.
+In Pre-LN, instability is further exacerbated when the statistical conditions assumed at initialization depart markedly from those observed later in training (§3.3).
+By contrast, Peri-LN alleviates this discrepancy, thereby improving training stability, and delivering additional performance gains.
+7.2 Precision Constraints Imposed by Pre-LN
+Both Pre-LN and Peri-LN architectures leave the main hidden state path unnormalized, so once large values arise in earlier layers, they persist through to later layers.
+Consequently, Pre-LN’s additive residual path might generates activations near or beyond the FP16 limit.
+To gauge how often these values exceed FP16 yet remain within BF16, we track the top-$100$ absolute hidden state values for $3.2$B-parameter Pre-LN and Peri-LN models.
+In Figure 11, the blue band for Pre-LN surpasses the FP16 maximum bound as early as $0.5$B training tokens whereas Peri-LN (red band) consistently remains below this threshold.
+This pattern, echoing Sun et al.
+(2024), highlights that choosing FP16 or BF16 is not just a hardware preference but is closely linked to how hidden state magnitudes evolve within the model.
+Earlier work on OPT (Zhang et al., 2022), which was pre-trained using FP16 precision, suggests that the training instabilities they observed were likely exacerbated by numerical overflows and gradient pathologies (Proposition 3.1) arising when activations exceeded the representable range of FP16.
+Figure 11: Evolution of extreme hidden state absolute magnitudes during training.
+Colored bands trace the range of the global top-$100$ absolute activations.
+Pre-LN (blue) quickly surpasses the FP16 representable maximum, while Peri-LN (red) stays below it.
 8 Conclusion
 We explore the placement of layer normalization within the Transformer architecture to better understand its role during training.
-By systematically comparing Post-LN, Pre-LN, and Peri-LN, we highlight their distinct impacts on stability, gradient propagation, and optimization dynamics.
+By systematically comparing Post-LN, Pre-LN, and newly termed Peri-LN, we highlight their distinct impacts on stability, final performance, and optimization dynamics.
 Our findings suggest that placing LN on module outputs in addition to the Pre-LN can help manage large activations while preserving beneficial gradient flow, thereby offering a promising balance for stable optimization.
@@
 We thank our colleague Jeongin Bae for inspiring the underlying motivation for this research.
-We are also grateful to Jung Hyun Lee, Seonghyeon Kim, and Seunghyun Seo for their valuable assistance during the early stages of our research discussions.
-Finally, we extend our gratitude to Gichang Lee, Lead of the Backbone Mission at NAVER Cloud, for his unwavering support in connecting us with key resources throughout this research.
+We are also grateful to Jung Hyun Lee, Seonghyeon Kim, and Seunghyun Seo for their valuable assistance during the early stages discussions.
+Finally, we extend our gratitude to Gichang Lee, Lead of the Backbone Mission at NAVER Cloud, for his unwavering support.
+This work was partly supported by Institute for Information & communications Technology Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT)(RS-2019-II190075, Artificial Intelligence Graduate School Support Program(KAIST), No.RS-2021-II212068, Artificial Intelligence Innovation Hub, No.
+RS-2024-00509279, Global AI Frontier Lab)
 Impact Statement
@@
 Given the growing industry focus on optimizing LLM deployment costs, our contributions are particularly relevant in the current AI research landscape.
-Another positive externality of improving the cost-effectiveness of large-scale training is the reduction in the environmental impact of AI research.
-Training state-of-the-art LLMs requires vast computational resources, leading to significant energy consumption and carbon emissions.
-By enhancing the efficiency of large-scale training, our approach helps mitigate these environmental costs, making AI research more sustainable.
-Lowering the energy footprint of training not only aligns with global efforts toward greener computing but also enables a wider range of institutions, including smaller research labs and academic groups, to participate in cutting-edge AI development without excessive resource constraints.
-This shift promotes a more inclusive and responsible AI ecosystem, ensuring that advancements in foundation models are both accessible and environmentally conscious.
+Improving the cost-effectiveness of large-scale training simultaneously lowers AI’s environmental footprint by reducing the vast energy consumption and carbon emissions inherent in state-of-the-art LLM development.
+This efficiency not only aligns with global sustainability goals but also enables smaller research labs and academic groups to pursue cutting-edge AI without prohibitive resource demands, fostering a more inclusive and responsible ecosystem.
 Appendix A Related Work
@@
 These findings illustrate the nuanced trade-offs of normalization placement and draw parallels to earlier CNN studies, where careful management of activation variance proved essential for stable deep learning.
+Ding et al.
+(2021) introduced Sandwich-LN in the vision domain for the first time, yet they paid little attention to the structural characteristics and differences that distinguish one LN placement from another.
+In language domain, several major open-source language models (e.g., Olmo2222 (OLMo et al., 2024), Gemma2222 (Rivière et al., 2024), and Gemma3333 (Team et al., 2025)) already employ a Peri-LN-like structure (see Section 3).
+Nevertheless, previous technical reports have not explained why this design might be advantageous compared with the more widely studied Pre-LN and Post-LN.
+By investigating Peri-LN in detail, we aim to highlight the structural benefits that appear to underlie its empirical success in these implementations.
 Gradient Propagation and Depth Scaling
@@
 While Post-LN and Pre-LN have received significant attention, we focus on a comparative analysis that includes Peri-LN, an alternative normalization placement that has thus far been underexplored but holds potential for enhancing training stability and model performance.
-Appendix B Detailed Experimental Setting
-In this section, we provide detailed configurations of both the pretraining and supervised fine-tuning to reproduce our results.
-B.1 Configurations on Pre-Training
-The common training settings are provided in Table 3.
-Embedding settings for the language models are listed in Table 4.
-For the model architecture, we primarily follow the Llama $3$ architecture (Dubey et al., 2024).
-In the MLP module, we use SwiGLU activations.
-Additional details regarding the model configurations are shown in Table 5.
-Note that embedding parameters are excluded from the model size.
-Unless otherwise noted, most training and model settings follow those of the DCLM experiments (Li et al., 2024a).
-During the pretraining stage, each model was trained under a controlled random seed.
-Table 3: Common configurations. LR Schedule denotes learning rate schedule.
-Global Batch Size Weight Decay Iterations Optimizer LR Schedule Warmup
-$256$ $0.033$ $14400$ Adam Cosine $10$%
-Table 4: Embedding configurations.
-Max Position Embeddings Position Embedding Type Untie-embeddings-and-output-weights
-$8192$ Rope $True$
-Table 5: Model configurations.
-Size $n_{layers}$ $n_{heads}$ $d_{model}$ $d_{head}$
-$400$M $24$ $8$ $1024$ $128$
-$1.5$B $24$ $16$ $2048$ $128$
-$3.2$B $32$ $16$ $2560$ $160$
-B.2 Configurations on Supervised Fine-Tuning
-To examine downstream task performance after instruction tuning, we employed a high-quality LIMA alignment training set consisting of $1,000$ samples (Zhou et al., 2023).
-Our supervised fine-tuning configuration was slightly modified from the original setup of LIMA: we fine-tuned the model for $15$ epochs with a batch size of $128$.
-The optimizer was Adam with an initial learning rate of $1$e-$5$ and a cosine learning rate schedule.
-We selected the best checkpoints for each model by evaluating on OpenBookQA (Mihaylov et al., 2018), CommonSenseQA (Talmor et al., 2019), and the NLI dataset in GLUE (Wang et al., 2018).
-Appendix C Additional Results on Pre-Training Study
-C.1 Post-Layer Normalization Architecture & Learning Rate Exploration
-In order to identify the optimal performance configuration for Post-LN within the experimental setup, we conducted a learning rate exploration as shown in Figure 8.
-Because the appropriate learning rate for Post-LN fell into a much lower range than those for Pre-LN and Peri-LN, we treated it separately.
-For each Post-LN setting, the best learning rate was determined as the one yielding the lowest final training loss, with the random seed held constant during this selection process.
-Figure 8: Learning rate explorations for Post-LN architectures.
-C.2 Best Performing Checkpoints Comparisons of Other Model Sizes
-As an extension of Section 4.2, we present below the results for additional model sizes that were omitted previously due to space constraints.
-In Figures 9, we compare the pre-training loss and the gradient norm curve at each LN strategy’s best-performing learning rate of $3.2$B and $1.5$B size models.
-(a) Training loss for $3.2$B
-(b) Gradient-norm for $3.2$B
-(c) Training loss for $1.5$B
-(d) Gradient-norm for $1.5$B
-Figure 9: Performance comparison of Post-LN, Pre-LN, and Peri-LN Transformers during pre-training for other two.
-C.3 Gradient Spikes during Pre-Training
-Although this observation is empirical, it remains sufficiently noteworthy to discuss.
-During extensive pre-training ablation experiments, we frequently observed gradient-norm spikes.
-Specifically, when examining how these spikes vary with different learning rates, we found that Pre-LN exhibited irregular spikes regardless of the learning rate setting, with a particularly high occurrence in the early stages of training.
-In contrast, such spikes were rarely observed under Peri-LN.
-Appendix D Proof of Theoretical Insight
-To support the claim that Peri-LN enhances the stability of training in such cases, we analyze the gradient norm in the final layer. For simplicity, we use RMSNorm and ReLU here.
-Proposition D.1.
+Appendix B Proposition 3.1 of Post-LN
+Proposition B.1.
+Post-LN (vanishing gradient). Consider the following sequence of operations:
+$a=\mathrm{MLP}(x),o=x+a,\tilde{o}=\mathrm{Norm}(o),$ (9)
+then
+$\left\lVert\frac{\partial\mathcal{L}(\tilde{o})}{\partial W_{i,j}^{(2)}}\right\rVert\;\leq\;\frac{4\,\gamma\,\sqrt{D}\,\|h\|}{\|x+a\|},$ (10)
+where $h:=\mathrm{ReLU}\left(xW^{(1)}+b^{(1)}\right)$.
+Since Post-LN normalizes the hidden state after each residual addition along the main path, the activation norm $|h|$ tends to remain relatively moderate.
+As a result, in Post-LN, the gradient scale is less sensitive to the magnitude of activations and more significantly influenced by model depth, as previously analyzed by Xiong et al.
+(2020) and Kedia et al.
+(2024), leading to vanishing gradients as depth increases.
+Appendix C Proof of Theoretical Insight
+To support the claim that Peri-LN enhances the stability of training in such cases, we analyze the gradient norm in the final layer.
+For simplicity, we use RMSNorm and ReLU here.
+Here, we assume that $\gamma$, the scaling parameter used in RMSNorm, is positive, and we empirically verified that it remains strictly positive across models of all sizes during training.
+Proposition C.1.
 Consider the following sequence of operations:
@@
 $\frac{\partial\mathcal{L}(o)}{\partial W_{i,j}^{(2)}}=h_{i}(\hat{p}_{j}-y_{j}).$ (20)
-Proposition D.2.
+Proposition C.2.
 Consider the following sequence of operations:
@@
 $\left\|\frac{\partial\mathcal{L}(o)}{\partial W_{i,j}^{(2)}}\right\|\leq 2\times 1\times\frac{2\gamma\sqrt{D}}{\|a\|}\times\|h\|=\frac{4\gamma\sqrt{D}\|h\|}{\|a\|}.$ (36)
-Proposition D.3.
+Proposition C.3.
 Consider the following sequence of operations:
@@
 The proof is analogous to the proof of the previous proposition. ∎
-Appendix E Additional Results on Growth of Hidden State
-In this section, we examine the $400$M- and $3.2$B-parameter models, which were omitted in Section 4.4.2 due to space constraints.
-As illustrated in Figures 10 and 11, these models exhibit the same overall trend.
+Appendix D Detailed Experimental Setting
+In this section, we provide detailed configurations of both the pretraining and supervised fine-tuning to reproduce our results.
+D.1 Configurations on Pre-Training
+The common training settings are provided in Table 2.
+Embedding settings for the language models are listed in Table 3.
+For the model architecture, we primarily follow the Llama $3$ architecture (Dubey et al., 2024).
+In the MLP module, we use SwiGLU activations.
+Additional details regarding the model configurations are shown in Table 4.
+Note that embedding parameters are excluded from the model size.
+Unless otherwise noted, most training and model settings follow those of the DCLM experiments (Li et al., 2024a).
+During the pretraining stage, each model was trained under a controlled random seed.
+Table 2: Common configurations. LR Schedule denotes learning rate schedule.
+Global Batch Size Weight Decay Iterations Optimizer LR Schedule Warmup Weight Initialization
+$256$ $0.033$ $14400$ Adam Cosine $10$% 0.02
+Table 3: Embedding configurations.
+Max Position Embeddings Position Embedding Type Untie-embeddings-and-output-weights
+$8192$ Rope $True$
+Table 4: Model configurations.
+Size $n_{layers}$ $n_{heads}$ $d_{model}$ $d_{head}$
+$400$M $24$ $8$ $1024$ $128$
+$1.5$B $24$ $16$ $2048$ $128$
+$3.2$B $32$ $16$ $2560$ $160$
+D.2 Configurations on Supervised Fine-Tuning
+To examine downstream task performance after instruction tuning, we employed a high-quality LIMA alignment training set consisting of $1,000$ samples (Zhou et al., 2023).
+Our supervised fine-tuning configuration was slightly modified from the original setup of LIMA: we fine-tuned the model for $15$ epochs with a batch size of $128$.
+The optimizer was Adam with an initial learning rate of $1$e-$5$ and a cosine learning rate schedule.
+We selected the best checkpoints for each model by evaluating on OpenBookQA (Mihaylov et al., 2018), CommonSenseQA (Talmor et al., 2019), and the NLI dataset in GLUE (Wang et al., 2018).
+Appendix E Additional Results on Pre-Training Study
+E.1 Post-Layer Normalization Architecture & Learning Rate Exploration
+In order to identify the optimal performance configuration for Post-LN within the experimental setup, we conducted a learning rate exploration as shown in Figure 12.
+Because the appropriate learning rate for Post-LN fell into a much lower range than those for Pre-LN and Peri-LN, we treated it separately.
+For each Post-LN setting, the best learning rate was determined as the one yielding the lowest final training loss, with the random seed held constant during this selection process.
+Figure 12: Learning rate explorations for Post-LN architectures.
+E.2 Best Performing Checkpoints Comparisons of Other Model Sizes
+As an extension of Section 4.2, we present below the results for additional model sizes that were omitted previously due to space constraints.
+In Figures 13, we compare the pre-training loss and the gradient norm curve at each LN strategy’s best-performing learning rate of $3.2$B and $1.5$B size models.
+(a) Training loss for $3.2$B
+(b) Gradient-norm for $3.2$B
+(c) Training loss for $1.5$B
+(d) Gradient-norm for $1.5$B
+Figure 13: Performance comparison of Post-LN, Pre-LN, and Peri-LN Transformers during pre-training for other two.
+Appendix F Additional Results on Growth of Hidden State
+In this section, we examine the $400$M- and $3.2$B-parameter models, which were omitted in Section 5.1 due to space constraints.
+As illustrated in Figures 14 and 15, these models exhibit the same overall trend.
 (a) Absolute magnitude growth
@@
 Figure 15: The forward growth patterns of hidden state for different architectures highlight the structural impact of normalization placement. $400$M size model.
-Appendix F Additional Experimental Results on Ablation Study
-F.1 Amount of Training Tokens
+Appendix G Additional Experimental Results on Ablation Study
+G.1 Amount of Training Tokens
 In order to investigate whether the learning behavior of each LN strategy varies with the number of training tokens, we conducted an additional round of learning-rate exploration for both the Pre-LN and Peri-LN architectures.
@@
 Figure 16: Learning rate explorations of Pre-& Peri-LN architecture with sequence length $2048$ configuration.
-F.2 Sequence Length
+G.2 Sequence Length
 In language models, the number of iterations per token is influenced by the sequence length, which in turn, along with the batch size, affects training statistics.
@@
 Figure 17: Learning rate explorations of Pre-& Peri-LN architecture with sequence length $2048$ configuration.
-F.3 Warm-up
+G.3 Warm-up
 Warmup is widely recognized to influence training stability.
@@
 Figure 18: Learning rate explorations of Pre-& Peri-LN architecture with warmup $30$% configuration.
-F.4 RMSNorm & LayerNorm
+G.4 RMSNorm & LayerNorm
 As illustrated in Figure 19, we conducted experiments in which RMSNorm and LayerNorm were interchanged.
@@
 Figure 19: LayerNorm vs. RMSNorm on Peri-LN architecture. $400$M size model.
-F.5 Embedding Layer Normalization of Peri-Layer Normalization Transformers
+G.5 Embedding Layer Normalization of Peri-Layer Normalization Transformers
 Motivated by Takase et al.
 (2023), we empirically explore the addition of embedding layer normalization to improve training stability and overall model performance in Transformer architectures.
-As illustrated in Figures 16, 17, and 18, incorporating Embedding LN in the Peri-LN architecture yields a slight improvement in pre-training loss.
+As illustrated in Figures 20, 21, and 22, incorporating Embedding LN in the Peri-LN architecture yields a slight improvement in pre-training loss.
 Furthermore, our empirical observations suggest that this effect becomes more pronounced in smaller models.
@@
 Figure 22: Loss and Gradient-norm curves comparing the presence and absence of Embedding LN in the Peri-LN architecture. $3.2$B size model.
-Appendix G Output-Layer Normalization with QK-Norm Architecture
+G.6 Ablation Study on Additional Normalization Layer Placement
+We additionally conduct further experiments on LN placements to compare different combinations (referred to as A, B, and C positions in Figure 2).
+We add configurations where LN is placed at both A + C (akin to combining Pre- and Post-LN), as well as only at B, to compare them with Peri-LN at final training loss under the controlled same training seed.
+We pre-train the $400$M-parameter Transformers on $30$B tokens each, using the same training configurations described in Section 4.
+As aligned with Xiong et al.
+(2020), our new results confirm that placing LN exclusively at C leads to training instability or suboptimal performance.
+In particular, the A + C configuration inherits characteristics of Post-LN (large gradient norm shifts), forcing the use of smaller learning rates and still resulting in lower overall performance than Peri-LN architecture.
+Table 5: Final training loss and additional normalization layer placement.
+$400$M A + C Post-LN B Peri-LN
+Final Training Loss $3.01$ $3.05$ Diverged $2.91$
+G.7 Peri-LN with QK-Norm
+While Peri-LN alone provides robust training dynamics, QK-Norm can still enhance performance.
+We conducted additional experiments that confirm combining Peri-LN with QK-Norm yields slight improvements in training loss.
+We pre-train the $1.5$B-parameter Transformers on $30$B tokens each, using the same training configurations described in Section 4.
+As shown in Table 6, adding QK-norm to Peri-LN indeed yielded better performance, consistent with Wortsman et al.
+(2024).
+In this experiment, the Peri-LN variant equipped with QK-norm used LayerNorm instead of RMSNorm.
+Table 6: Peri-LN with QK-Norm.
+$1.5$B Peri-LN +QK-Norm (Wortsman et al., 2024)
+Final Training Loss $2.722$ $2.711$
+G.8 Weight Decay and Weight Initialization
+G.8.1 Common Settings
+We pre-train the $400$M-parameter Transformers on $30$B tokens each under the controlled same training seed.
+We measure the training loss and averaged benchmark score for these experiments under the same evaluation settings used in Table 1 of the paper.
+Other configurations follow those outlined in Section 4.
+For the variance growth experiments in Figure 10, we adopt the same settings as in Section 5.1, except that we use $100$ samples for the forward-pass statistics.
+G.8.2 Weight Decay
+We conduct additional studies for various weight decay condition for both Pre-LN and Peri-LN architectures.
+As shown in the Table 7, Peri-LN continues to offer better performance than Pre-LN under the same settings.
+We provide per-run results as below:
+Table 7: Effect of weight decay on $400$M-parameter Pre-LN and Peri-LN Transformers: Final training loss and averaged benchmark score.
+$400$M Weight Decay Coefficient 00 $0.0033$ $0.033$ $0.33$
+Final Training Loss Pre-LN $3.03$ $3.03$ $3.03$ $3$
+Peri-LN $2.94$ $2.94$ $2.93$ $2.90$
+Averaged Benchmark Score Pre-LN $49.26$ $49.18$ $49.01$ $49.51$
+Peri-LN $51.41$ $51.14$ $50.68$ $52.13$
+(a) Pre-training loss curve
+(b) Gradient-norm curve
+Figure 23: Comparison of pre-training loss and gradient norms for Pre-LN and Peri-LN architectures with the weight decay coefficient fixed at 00, while all other hyperparameters are held constant.
+(a) Pre-training loss curve
+(b) Gradient-norm curve
+Figure 24: Comparison of pre-training loss and gradient norms for Pre-LN and Peri-LN architectures with the weight decay coefficient fixed at $0.0033$, while all other hyperparameters are held constant.
+(a) Pre-training loss curve
+(b) Gradient-norm curve
+Figure 25: Comparison of pre-training loss and gradient norms for Pre-LN and Peri-LN architectures with the weight decay coefficient fixed at $0.033$, while all other hyperparameters are held constant.
+(a) Pre-training loss curve
+(b) Gradient-norm curve
+Figure 26: Comparison of pre-training loss and gradient norms for Pre-LN and Peri-LN architectures with the weight decay coefficient fixed at $0.33$, while all other hyperparameters are held constant.
+G.8.3 Weight Initialization
+We run an additional ablation on weight-initialization schemes.
+For both Pre-LN and Peri-LN models, we first adopt Xavier initialization (Glorot & Bengio, 2010) and then compare it with He initialization ($2/d$) (He et al., 2015), LeCun initialization ($1/d$), and two scaled variants, $10/d$ and $1/(10d)$.
+As Table 8 shows, Xavier initialization yields the strongest overall performance, improving on the configurations used in our earlier experiments.
+Crucially, our central finding remains intact: hidden-state variance sharply grows in Pre-LN Transformers but stays bounded in Peri-LN Transformers.
+Table 9 confirms the same pattern across all initialization scales, and detailed per-run results appear below.
+Table 8: Xavier initialization (Glorot & Bengio, 2010) yields better performance compared to our previous weight initialization configurations.
+400M Architecture Baseline($0.02$) Xavier Initialization
+Loss Pre-LN 3.03 2.95
+Peri-LN 2.93 2.91
+Avg. Pre-LN 49.01 51.25
+Peri-LN 50.68 52.04
+Table 9: Effect of weight-initialization variance on final pre-training loss for $400$M-parameter Pre-LN and Peri-LN Transformers.
+400M Initialization Variance 10/d He (2/d) LeCun (1/d) 1/(10d) Baseline ($0.02$)
+Loss Pre-LN 4.526 2.965 3.005 3.012 3.035
+Peri-LN 3.027 2.929 2.915 2.902 2.916
+(a) Pre-training loss curve
+(b) Gradient-norm curve
+Figure 27: Comparison of pre-training loss and gradient norms for Pre-LN and Peri-LN architectures with the weight initialization variance set to $10/d$, while all other hyperparameters are held constant.
+(a) Pre-training loss curve
+(b) Gradient-norm curve
+Figure 28: Comparison of pre-training loss and gradient norms for Pre-LN and Peri-LN architectures with the weight initialization variance set to $2/d$ (He init (He et al., 2015)), while all other hyperparameter are held constant.
+(a) Pre-training loss curve
+(b) Gradient-norm curve
+Figure 29: Comparison of pre-training loss and gradient norms for Pre-LN and Peri-LN architectures with the weight initialization variance set to $1/d$ (LeCun init), while all other hyperparameter are held constant.
+(a) Pre-training loss curve
+(b) Gradient-norm curve
+Figure 30: Comparison of pre-training loss and gradient norms for Pre-LN and Peri-LN architectures with the weight initialization variance set to $1/(10d)$, while all other hyperparameter are held constant.
+Appendix H Output-Layer Normalization with QK-Norm Architecture
 Query and Key layer-normalization (QK-Norm) has been widely studied in modern Transformer architectures (Wortsman et al., 2024; Zhai et al., 2023; OLMo et al., 2024).
@@
 In our view, QK-Norm and Pre-LN (placed at $A^{2}$ and $A$ respectively in Figure 31) serve the same role but differ in certain details.
-As shown in Figures 20, 21, and 22, the two architectures exhibit comparable performance overall in terms of both training loss and stability..
+As shown in Figures 32, 33, and 34, the two architectures exhibit comparable performance overall in terms of both training loss and stability..
 However, Peri-LN provides a slight performance advantage over the OLMo2222-style Peri-LN in the $400$M- and $1$B-parameter models.
@@
 $3.2$B size model.
-Appendix H Additional Details on Evaluation
-H.1 Detailed Configurations
+Appendix I Training the Transformer using Stochastic Gradient Descent
+Using Stochastic Gradient Descent (SGD) for training Transformers is not a common practice.
+As Zhang et al.
+(2024) point out, Transformer-based models tend to perform worse with SGD than with Adam by a considerable margin.
+One reason is that SGD struggles to handle the heterogeneity across different blocks.
+Although these aspects are certainly intriguing and warrant further investigation, they lie beyond the scope of our current work, as Zhang et al.
+(2024) also note.
+Nonetheless, for someone who might wonder to know, we conduct additional experiments using SGD.
+We are searching for U-shaped patterns during the learning rate exploration for both Pre-LN & Peri-LN as shown in the Figure 35.
+We observe that: (1) SGD performs worse than Adam, consistent with findings reported in (Zhang et al., 2024); and (2) Peri-LN demonstrates better performance than Pre-LN.
+In these results, we use $400$ M-parameter Transformers and apply the same configurations as in the main experiments (Section 4.1).
+We provide detailed training curves in Figures 363738.
+Figure 35: Learning Rate Exploration of Pre-& Peri-LN architecture trained with SGD optimizer. The individual training-loss and gradient-norm curves appear in Figures 36, 37, and 38.
+(a) Pre-training loss curve
+(b) Gradient-norm curve
+Figure 36: Using SGD with learning rate $35e-$ 3.
+(a) Pre-training loss curve
+(b) Gradient-norm curve
+Figure 37: Using SGD with learning rate $33e-$ 3.
+(a) Pre-training loss curve
+(b) Gradient-norm curve
+Figure 38: Using SGD with learning rate $31e-$ 3.
+Appendix J Additional Details on Evaluation
+J.1 Detailed Configurations
 We utilized the Language Model Evaluation Harness [^6] https://github.com/EleutherAI/lm-evaluation-harnesswith the HuggingFace Transformers library (Gao et al., 2024; Wolf et al., 2020) to assess overall performance.
@@
 When calculating the evaluation score, diverged checkpoints were excluded.
-H.2 Detailed Results on Benchmark Evaluations
+J.2 Detailed Results on Benchmark Evaluations
 In this section, we present the evaluation results for each model trained with five different training seeds.
 We exclude any diverged scores and average the remaining values, which are then reported in Table 1 in the main text.
-H.2.1 Pre-Training
+J.2.1 Pre-Training
 Table 10: Detailed results on pre-training the Peri-LN architecture. These results are averaged to produce the values reported in Table 1. SEED denotes pre-training seed.
@@
 5 0.4663 0.3340 0.6839 0.3577 0.5043
-H.2.2 Supervised Fine-Tuning
+J.2.2 Supervised Fine-Tuning
 Table 13: Detailed results on SFT with Peri-LN architecture. These results are averaged to produce the values reported in Table 1. SEED denotes pre-training seed.
@@
 5 0.5341 0.3677 0.6942 0.3976 0.5012
+Appendix K Additional Discussions of Precision Constraints Imposed by Pre-LN
+This section provide additional discussions of Section 7.2.
+Given that both Pre-LN and Peri-LN exhibit a structural property whereby large values do not readily disappear once they arise, it is important to monitor the occurrence of these extreme activations.
+Pre-LN’s additive residual path often produces activations that approach, and occasionally exceed, the FP16 (float16) representable maximum [^7] https://en.wikipedia.org/wiki/Bfloat16_floating-point_format.
+To quantify how often these values would overflow FP16 yet remain within the BF16 (bfloat16) range [^8] https://en.wikipedia.org/wiki/Half-precision_floating-point_format, we measure the 100 largest absolute hidden-state values (the global top-$100$ tokens) for each Pre-LN and Peri-LN $3.2$B-parameter Transformer.
+The shaded region indicates the range of these global top-$100$ tokens.
+The blue curve (with shaded band) represents a Pre-LN model, the red curve a Peri-LN model, and the dashed orange line denotes the FP16 representable maximum.
+As shown in Figure 11, consistent with the observations of Sun et al.
+(2024), activations in the Pre-LN model routinely exceed the FP16 bound from the very first $0.5$B training tokens, with the overshoot becoming more pronounced in deeper layers—an indication of growing numerical instability.
+By contrast, Peri-LN consistently maintains a comfortable margin below the FP16 limit throughout training.
+This finding underscores that the choice between FP16 and BF16 is not merely a hardware preference; it is tightly coupled to how hidden-state magnitudes evolve within the network.
+Since NVIDIA V100 GPUs do not support BF16, these results imply that training and inference with Pre-LN models on such hardware are inherently disadvantaged.
+Moreover, from the standpoint of large-language-model quantization (Dettmers et al., 2022; Lee et al., 2023; Kim et al., 2023; Heo et al., 2024; Lee et al., 2025), the prevalence of massive activations in Pre-LN can severely disrupt outlier-aware compression algorithms.
+When aggressive low-precision compression is the goal, the Pre-LN architecture’s tendency to generate extreme hidden state values therefore constitutes a particularly challenging obstacle.
+Meta AI’s publicly released OPT training logbooks and chronicles document repeated episodes of gradient divergence and loss spikes encountered while pre-training entirely in FP16 precision  [^9] https://github.com/facebookresearch/metaseq/blob/main/projects/OPT/chronicles/10_percent_update.md.
+Since FP16 saturates at an absolute value of $65,504$, any hidden state activation that exceeds this threshold silently overflows, corrupting the forward pass and, through back-propagation, inducing erratic gradients.
+Earlier analyses of OPT (Zhang et al., 2022) therefore ascribe much of the observed instability to numerical overflow, a view formalized in our Proposition 3.1, which shows how such out-of-range activations propagate into severe gradient pathologies.
+These external reports provide further corroboration that architectures prone to generating large-magnitude activations—as Pre-LN does—require either a wider numerical format (e.g., BF16) or explicit magnitude-regularization to ensure stable large-scale training.
+Appendix L Additional Discussions of Hidden State Representations
+As shown in Figure 39(a), Post-LN exhibits smaller angular distances due to the LN being located on the main path, whereas Pre-LN and Peri-LN begin with very similar states.
+As shown in Figure 39(c), at the end of training, Pre-LN tends to produce more redundant hidden state representations compared to the others.
+This phenomenon may stem from Pre-LN’s repeated residual additions, which amplify certain representations over others.
+We use $30$B tokens trained $400$M size model in this experiments.
+For dataset, we utilize $256$ random samples of RedPajama-Data-$1$T (Computer, 2023) for this results.
+To investigate further, we focus on module-output normalization, the primary factor distinguishing Pre-LN from Peri-LN.
+As shown in Figure 39(b), the learnable scale starts around $1$ in the early stages of training and gradually changes with increasing depth.
+Because Peri-LN preserves the identity path, it appears to adjust the scale of the module output accordingly.
+This suggests that the exponential growth of the main path’s magnitude in Pre-LN diminishes the relative contribution of individual modules, resulting in more redundant hidden representations.
+Figure 39(d) shows that fixing the learnable scale of Peri-LN’s module output LN at $1$ causes the main path contribution to decrease in deeper layers.
+This finding confirms the role of module output normalization in controlling hidden state redundancy.
+(a) At initialization
+(b) Learnable scale $\gamma$ in Output-LN
+(c) After $30$B tokens training
+(d) Frozen $\gamma$ in Output-LN
+Figure 39: Angular distance of hidden state is presented in Figure  39(a)39(c), and 39(d).
+In Figure 39(b), we monitor $\gamma$ of every Output-LN in Peri-LN during training.
+We use $30$B tokens trained $400$M size model in this experiments.